### v1

# Atomic Fact Decomposition Helps Attributed Question Answering

Abstract.
Attributed Question Answering (AQA) aims to provide both a trustworthy answer and a reliable attribution report for a given question. Retrieval is a widely adopted approach, including two general paradigms: Retrieval-Then-Read (RTR) and post-hoc retrieval. Recently, Large Language Models (LLMs) have shown remarkable proficiency, prompting growing interest in AQA among researchers. However, RTR-based AQA often suffers from irrelevant knowledge and rapidly changing information, even when LLMs are adopted, while post-hoc retrieval-based AQA struggles with comprehending long-form answers with complex logic, and precisely identifying the content needing revision and preserving the original intent. To tackle these problems, this paper proposes an Atomic fact decomposition-based Retrieval and Editing (ARE) framework, which decomposes the generated long-form answers into molecular clauses and atomic facts by the instruction-tuned LLMs. Notably, the instruction-tuned LLMs are fine-tuned using a well-constructed dataset, generated from large scale Knowledge Graphs (KGs). This process involves extracting one-hop neighbors from a given set of entities and transforming the result into coherent long-form text. Subsequently, ARE leverages a search engine to retrieve evidences related to atomic facts, inputting these evidences into an LLM-based verifier to determine whether the facts require expansion for re-retrieval or editing. Furthermore, the edited facts are backtracked into the original answer, with evidence aggregated based on the relationship between molecular clauses and atomic facts. Extensive evaluations demonstrate the superior performance of our proposed method over the state-of-the-arts on several datasets, with an additionally proposed new metric $Attr_{p}$ for evaluating the precision of evidence attribution.
Attributed Question Answer, Information Retrieval, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; April 28 - May 2; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 184ccs: Information systems Information retrievalccs: Computing methodologies Natural language generation
1. Introduction
Large Language Models (LLMs), pre-trained on large-scale text corpora (10387715, ), have demonstrated remarkable capabilities in natural language understanding and generation tasks (zhuang2023toolqa, ; wang2024large, ). However, despite their impressive performance, they often face significant challenges in real-world applications due to issues like lack of interpretability and indecisiveness (huang2023survey, ; peskoff-stewart-2023-credible, ). These limitations undermine the reliability and trustworthiness of LLMs, particularly in contexts where transparency and accuracy are essential. To address these issues, Attributed Question Answering (AQA) (li2023survey, ) has emerged as a solution, focusing on linking their generated answers to specific sources of evidence.
Figure 1. The motivation of previous methods and our proposed ARE. Previous methods directly use LLMs to generate sub-questions for the long-form answer and edit the whole answer. Our proposed ARE leverages instruction-tuned LLMs to achieve molecular-to-atomic stage, and backtrack the edited atomic facts to achieve atomic-to-molecular stage.
Depending on the timing of retrieval, previous research can be divided into two categories: Retrieval-then-Read (RTR)  (gao2023enabling, ), and Post-hoc Retrieval (bohnet2022attributed, ; gao2023rarr, ; chen2023purr, ). RTR delivers relevant answers by retrieving documents based on the query, providing detailed context that can enhance the richness of the response. However, LLMs often link content to irrelevant or incorrect sources and lack an effective mechanism for subsequent correction. In contrast, post-hoc retrieval typically retrieves specific external information for initially generated long-form answers, effectively reducing linking errors and enabling targeted revisions, which enhances flexibility and robustness. (dhuliawala2023chainofverification, ; kim2024re, ).
Despite the flexibility post-hoc retrieval offers in verifying each fact and correcting it in long-form answers, it introduces several significant challenges. Firstly, existing methods (chen2023purr, ; kim2024re, ) prompt LLMs to directly generate a series of sub-questions for retrieval based on the long-form answer (As shown in Figure 1 (I)). However, the complex logical structures and vast amounts of information typically found in long-form answers make it difficult for these methods to comprehensively capture all necessary facts. As a result, the generated sub-questions often fail to align with the core content of the answer, leading to irrelevant questions and ultimately hindering the retrieval of accurate supporting evidence. tribute methods always overlooking the broader context, which can lead to inaccurate or incomplete attribution.
Moreover, existing methods (gao2023rarr, ; wu2024updating, ) that directly edit long-form answers holistically often face difficulty in precisely locating specific content for revision, making it challenging to preserve the original intent while ensuring consistency between the answer and the attribution. This limitation arises from either insufficient editing, where critical details remain inadequately revised, or excessive editing, which distorts the original intent or introduces unnecessary changes, which disrupts the coherence of the answer. Thus, effectively balancing the challenges of insufficient and excessive editing has become a critical issue.
To address the above limitations, we propose a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for AQA (As shown is Figure 1 (II)). Specifically, ARE first prompts LLMs to generate long-form answers for the given question. Subsequently, ARE utilizes an instruction-tuned LLM, trained on a well-constructed fact decomposition dataset, to decompose the long-form answers into molecular clauses and then into atomic facts. These atomic facts are then used to search the evidences from a search engine.
Additionally, ARE employs an evidence verifier to classify the relationships between the evidence and the atomic facts into three categories: 1) supportive, 2) editing required, and 3) irrelevant. When an atomic fact requires editing or is irrelevant to the retrieved evidence, ARE uses LLMs to process the atomic fact accordingly: if the relationship is editing required, ARE revises the fact; if the relationship is irrelevant, ARE expands the atomic fact for re-retrieval and verification. This process is repeated until the verification result shows that the relationship is “supportive” or the maximum number of iterations is reached.
Finally, ARE backtracks the edited atomic facts to their original positions within the molecular clauses, forming the final revised answer while preserving the original intent. Meanwhile, the evidences are aggregated based on the relationships between the molecular clauses and atomic facts to generate the attribution report. Meanwhile, ARE introduces a more comprehensive evaluation metric $Attr_{p}$, which not only accurately assesses the precision of retrieved evidence, but also emphasizes the completeness of the evidence. Our contributions are summarized as follows:
  • This paper proposes a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for the post-hoc retrieval-based AQA task, which performs editing and verification of long-form answers at both molecular and atomic levels;
  • This paper introduces an innovative instruction-tuned fact decomposition LLM, which is fine-tuned on a carefully constructed molecular-to-atomic fact decomposition dataset;
  • This paper proposes a traceable editing method that performs atomic fact editing and evidence retrieval at the atomic level, while ensuring consistency between atomic facts and evidences, and preserving the original intent at the molecular level;
  • This paper designs an evaluation metric $Attr_{p}$ to accurately assess the precision and completeness of retrieved evidences, mitigating the proportion of invalid retrieved evidences.
Figure 2. An example of the ARE process. In Molecular-to-Atomic stage, an instruction-tuned fact decomposition LLM decomposes the long-form answer $X$ into molecular clauses and atomic facts. Atomic facts are then used to retrieve evidences. In Atomic-to-Molecular stage, a Verifier assesses the relationship between the evidence and facts, resulting in three states: 1. supportive, requiring no further action; 2. editing required, necessitating revision to the atomic facts; and 3. irrelevant, requiring a new evidence retrieval. Finally, the atomic facts are backtracked to the original position of $X$ to generate the revised answer $X^{{}^{\prime}}$. The Attribution Report consists of all evidences $\{E_{1},...,E_{m}\}$, where $MF_{i}^{{}^{\prime}}$ and $e_{ij}^{{}^{\prime}}$ are the edited contents.
2. Related Work
2.1. Attributed Question Answering
From the perspective of retrieval timing, two notable trends have recently emerged: (1) Retrieval-then-Read (RTR) and (2) Post-hoc Retrieval. Trivedi et al. (trivedi2022interleaving, ) introduce IRCoT, a method that interleaves retrieval with steps in a Chain of Thought (CoT). This approach not only guides the retrieval process using CoT but also utilizes the retrieved results to enhance the CoT.
Muller et al. (muller-etal-2023-evaluating, ) investigate attribution in cross-lingual question answering (QA). ALCE (gao2023enabling, ) employs various methods to integrate retrieved documents into large language models (LLMs) for answer generation. Li et al. (li2023llatrieval, ) introduce a progressive selection of evidence using LLMs with a classification-based prompting template. Bohnet et al. (bohnet2022attributed, ) conduct an extensive evaluation of LLM attributions, finding that while the Retrieval-then-Read (RTR) approach performs well, it necessitates the comprehensive use of a traditional training set, thereby highlighting the potential of post-hoc retrieval. RARR (gao2023rarr, ) is the first framework to implement question decomposition-based retrieval followed by revision. Building on RARR, PURR presents an end-to-end editor for text revision (chen2023purr, ). Kang et al. (kang2023mitigating, ) propose a combination of RTR and post-hoc retrieval strategies. The limitations of question decomposition were discussed in the Introduction. Our approach addresses these limitations through molecular-to-atomic fact decomposition and atomic-to-molecular editing, resulting in significant performance improvements.
2.2. Fact Decomposition
The technology of decomposition has been shown to effectively address complex questions, particularly in various reasoning tasks (lin2023decomposing, ; Ye_2023, ; zhang2023reasoning, ) and claim verification tasks (lin-etal-2023-decomposing, ; chen2023complex, ; kim2023kggpt, ; wang2023hallucination, ; kamoi-etal-2023-wice, ). While fact decomposition can accurately generate sub-facts that represent the original answer, it struggles to aggregate evidence for sentences based on the decomposition results. This limitation becomes evident when retrieving evidence by fact to support each sentence that contains more than one fact.
2.3. Hallucination Detection and Revision
Hallucination detection (chen2023hallucination, ; wang2023hallucination, ; kadavath2022language, ; mndler2023selfcontradictory, ; mishra2024fine, ) is a challenging yet essential task for improving the reliability of large language models (LLMs) in real-world scenarios. Similarly, Zheng et al. (ZLLP2024, ) introduced TrustScore, the first effective evaluation metric designed to assess the trustworthiness of LLM responses in a reference-free context.
To tackle hallucinations, numerous model editing methods have been developed that do not involve updating the parameters of large language models (LLMs) (zheng-etal-2023-edit, ; gao2023rarr, ; HLTW+2023, ; dhuliawala2023chainofverification, ; wu2024updating, ; song2024knowledge, ). In contrast to our approach, most of these methods focus exclusively on either hallucination detection or revision.
3. Preliminary
Attributed Question Answering (AQA) is a task which provides both a trustworthy answer and a reliable attribution report for a given question. Formally, given a question $q$ and a corpus of text passages $D$, the process can be defined as follows:
$X,\ A\longleftarrow\mathcal{M}_{\text{AQA}}(q,\ D),$
where $X$ represents the long-form answer, $A$ is an attribution report consisting of of a collection of evidence, and $\mathcal{M}_{\text{AQA}}$ is a model used for the AQA task.
Post-hoc Retrieval-based AQA aims to enhance the reliability of LLM-generated long-form answers by further incorporating evidence retrieval, fact verification, and factual editing. A well-designed framework should maximize the attribution score while minimizing changes to the original intent.
Symbol definitions. The key symbols used in our framework are depicted as follows: $MF_{i}$ represents a molecular clause within the long-form answer $X$; $AF_{ij}$ denotes the atomic fact of $MF_{i}$; $e_{ij}$ corresponds to the evidence for each atomic fact $AF_{ij}$; $E_{i}$ represents the aggregated evidence for the molecular clause $MF_{i}$, compiled from multiple evidence $e_{ij}$.
Entropy measures the uncertainty or information content associated with random variables. The higher the entropy, the more uncertain or random the information is (ash2012information, ). It can be calculated as:
$H(X)=-\sum_{i=1}^{n}P(x_{i})\log P(x_{i}),$
where $P(x_{i})$ is the probability of $x_{i}$, which can be defined as:
$P(x_{i})=P(t_{1})\cdot P(t_{2}\mid t_{1})\cdot\dots\cdot P(t_{i}\mid t_{1},t_{2},\dots,t_{i-1}),$
where $t_{i}$ is $i$-th token in $x_{i}$.
Hypothesis. Atomic facts exhibit lower entropy than sub-questions, leading to improved evidence retrieval. As demonstrated by Passalis et al. (7439840, ) that minimizing entropy can enhance performance in retrieval tasks. Long-form answers, characterized by greater complexity, have higher entropy, which makes it more challenging to retrieve relevant evidences. Traditional post-hoc retrieval based AQA methods use LLMs to generate sub-questions directly, which may contain multiple meanings and uncertainties. Thus, these sub-questions often have high entropy, leading to retrieving ambiguous or irrelevant evidence (As shown in Figure 1 (I)). In contrast, fact decomposition preserves higher certainty and lowers entropy, thereby enhancing evidence retrieval (ash2012information, ). The experimental results in Section 5.5 verify our hypothesis by demonstrating that atomic-fact based retrieval outperforms sub-question based retrieval.
Figure 3. The process of dataset construction and instruction tuning. We first extract the neighbors of the corresponding entity from Wikidata, and then utilize an LLM to convert each triple into a single atomic fact. Subsequently, all facts are transformed into a text comprising multiple molecular clauses by the LLM. The atomic facts are then used to construct each sentence. After two prompts, we obtain the dataset for instruction-tuning with LoRA.
4. Approach
In this section, a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework is proposed for AQA task. Specifically, our ARE framework contains three stages: 1) Molecular-to-Atomic Fact Decomposition generally decomposes the generated long-form answer into molecular clauses and atomic facts by an instruction-tuned LLM; 2) Atomic Fact based Evidence Retrieval employs a search engine to retrieve evidence and selects the most similar evidence using a reranking model; 3) LLM-based Evidence Verifing and Editing guides LLMs through a verified result to either expand atomic facts for further re-retrieval or edit atomic facts with retrieved evidence.
4.1. Molecular-to-Atomic Fact Decomposition
This stage introduces the process that prompts LLMs to generate a long-form answer, and utilizes the instruction-tuned LLMs fine-tuned on a carefully constructed dataset to decompose the long-form answer into molecular clauses and atomic facts.
4.1.1. Long-form answer generation
LLMs are trained on the large-scale corpora which endow them with powerful generative capabilities. Long-form answers are generated by feeding few-shot prompting template and a specific question to LLMs. Formally, the whole process can be represented as:
Question: Player with most Super Bowl rings? Long-form Answer $X$: The player with the most Super Bowl rings is Tom Brady. Tom Brady is an American football quarterback who has won six Super Bowl championships.
4.1.2. Fact decomposition
Fact decomposition aims to decompose long-form answers characterized by complex logical structures into molecular clauses and multiple atomic facts. Specifically, our ARE framework constructs a molecular-to-atomic fact decomposition dataset for instruction-tuning LLM, and this process enhances LLM’s ability to effectively manage the intricacies of long-form answers.
Dataset Construction. To fine-tune the molecule-to-atomic fact decomposition LLM, ARE constructs an instruct-tuning dataset that converts triples into texts containing molecule clauses and atomic facts derived from Knowledge Graphs (KGs) using LLMs. As depicted in Figure 3, the dataset construction process is divided into two steps: data collection and data processing with LLMs.
Step1: ARE first randomly selects the entities from the Knowledge Graph Question Answering (KGQA) dataset (gu2021beyond, ) and then extracts the triplets of given entities from Wikidata [^1] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=. As some properties or objects in the triplets may be useless (e.g., JPEG, code, and ISSN), they are thus further filtered by heuristic rules in the extracting process. Details of the rules can be found in Appendix C.4.
Step2: ARE utilizes LLMs to randomly transform several triplets into atomic facts $\{AF_{i1},...,AF_{ik}\}$. The atomic facts are fed into LLMs to generate a response in the JSON format: {“Generated content”: $S$, “Molecular clauses”: $MF_{i}$, “Atomic facts”: $AF_{ij}$}, where $S$ is text consisting of $n$ molecular clauses, $MF_{i}$ is $i$-th molecular clause, $i\in\{1,...,m\}$, and $AF_{ij}$ is the $j$-th atomic fact of the $MF_{i}$, $j\in\{1,...,k\}$.
Through the above process, an instruction-tuning dataset can be obtained, which contains 7138 samples for training and 1000 for evaluation. An example data can be found in Appendix A.2.
Instruction-tuning the Molecule-to-Atomic LLM. For the fine-tuning stage, ARE employs the Low-Rank Adaptation (LoRA) technique (hu2022lora, ) for instruction-tuning the Llama3-8B-Instruct LLM to reduce computational complexity and time consumption. Specifically, LoRA works by freezing parameters $\theta_{0}$ of the pre-trained model, while adding trainable parameters $\Delta\theta_{0}$ that can be expressed as the product of two low-rank matrices:
(1) $\Delta\theta_{0}=\mathbf{BA},$
where $\mathbf{B}\in\mathbb{R}^{d\times r}$, $\mathbf{A}\in\mathbb{R}^{r\times k}$, $r\ll\mathrm{min}(d,k)$.
After obtaining the instruction-tuned Molecule-to-Atomic LLM, ARE decomposes the long-form answer “$X$: The player with the most Super Bowl rings is Tom Brady. Tom Brady is an American football quarterback who has won six Super Bowl championships.” into molecular clauses “$MF_{1}$: The player with the most Super Bowl rings is Tom Brady.” and “$MF_{2}$: Tom Brady is an American football quarterback who has won six Super Bowl championships.”, $MF_{1}$ only has one atomic fact $AF_{11}$, $MF_{2}$ can be further decomposed into two atomic facts “$AF_{21}$: Tom Brady is an American football quarterback.” and “$AF_{22}$: Tom Brady has won six Super Bowl championships.”.
4.2. Atomic Fact based Evidence Retrieval
Atomic fact based evidence retrieval aims to employ search engines [^2] We select Google Search as knowledge source, accessible via https://customsearch.googleapis.com/customsearch/v1 to search evidences that support atomic facts. To be specific, ARE utilizes a pre-trained Sentence-Bert model to generate embeddings for atomic facts and their corresponding evidences, and subsequently re-ranks evidences based on similarity to identify the most relevant one.
ARE leverages search engines to search evidence from each atomic fact $AF_{ij}$, while uses the pre-trained Sentence-Bert model [^3] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2 to embed both the searched evidences $e_{c}$ and the atomic fact $AF_{ij}$ into a common vector space:
(2) $\mathbf{e}_{c}^{z},\text{{AF}}_{ij}=\textrm{Sentence-Bert}(e_{c}^{z},AF_{ij}),$
where $e_{c}^{z},z\in\{1,...,o\}$ is the $z$-th searched evidence in $e_{c}$. ARE calculates the relevance score $R$($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) through the cosine similarity function:
(3) $R(\mathbf{e}_{c}^{z},\text{{AF}}_{ij})=\frac{\mathbf{e}_{c}^{z}\cdot\text{{AF}}_{ij}}{\parallel\mathbf{e}_{c}^{z}\parallel\parallel\text{{AF}}_{ij}\parallel},$
[…]
4.3. LLM-based Evidence Verifing and Editing
Due to the fact that hallucinated knowledge probably exists in the long-form answers generated by LLMs, ARE utilizes the evidence verifing LLM to assess whether atomic facts need to be edited or re-retrieved through comparing with the retrieved evidence. The evidence verifier takes as input the atomic fact and its corresponding evidence, and outputs different statuses, including:
(4) $\textrm{EV}(AF_{ij},e_{ij})=\left\{\begin{matrix}1,&\textit{supportive}\\ 2,&\textit{editing required}\\ 3,&\textit{irrelevant},\end{matrix}\right.$
where $\text{EV}(\cdot,\cdot)$ denotes the evidence verifier, which determines the relationships between the atomic fact and evidence. Specifically, $1$ represents supportive, requiring no further action; $2$ means editing required, necessitating the revision of atomic facts; $3$ indicates irrelevant, which requires a new evidence retrieval. The prompt template of the evidence verifier can be found in Appendix A.3.
Figure 2 illustrates the process. The status between $AF_{11}$ and $e_{11}$ is classified as supportive, $AF_{21}$ retrieves irrelevant evidence $e_{21}$, thus requiring re-retrieval. $AF_{22}$ needs to be edited based on evidence $e_{22}$. For irrelevant, ARE utilizes the fact expanding LLM to expand the atomic fact $AF_{21}$. The expanded fact contains more complete information, enabling more effective retrieval of relevant evidence. For example, ”Tom Brady” is expanded to ”Thomas Edward Patrick Brady Jr.”. The newly retrieved evidence will be re-verified by the evidence verifier, and this process will be repeated. Details of the prompt template and the effectiveness of fact expansion from the perspective of the information theory can be found in Appendix A.3 and Appendix B.
For editing required, ARE designs prompts for LLMs based on in-context learning and chain-of-thought prompting techniques  (zheng-etal-2023-edit, ; gao2023rarr, ), guiding the LLMs to revise the atomic fact $AF_{22}$ using the retrieved evidence $e_{22}$. Since the edits are made at the atomic fact level rather than revising the entire answer $X$, ARE can precisely adjust the necessary details, thus minimizing unnecessary modification. The detailed editing instructions are shown in Appendix A.4.
The edited atomic facts are re-verified and backtracked to their original positions within the molecular clauses, forming the final revised long-form answer $X^{\prime}$ along with the other molecular clauses that do not require editing. To further obtain the attribution report $A$, all evidences $e_{ij}$ are aggregated into a sequence $E_{i}$ to support the $MF_{i}$. Since each $AF_{ij}$ is derived from the decomposition of $MF_{i}$, there may be overlaps among the evidence $e_{ij}$. To address this issue, duplicate snippets are removed. Ultimately, $A$ is compiled as $\{E_{1},...,E_{i},...,E_{m}\}$.
[…]
Methods GPT-3.5 Llama3-70B Llama2-70B
[…]
NQ
DRQA 0.424 0.647 - - - 0.382 0.620 - - - 0.483 0.522 - - - - -
EFEC 0.598 0.042 0.762 0.080 0.670 0.490 0.058 0.717 0.107 0.582 0.357 0.058 0.719 0.107 0.477 0.098 0.576
CCVER 0.624 0.066 0.928 0.123 0.747 0.597 0.071 0.921 0.132 0.724 0.423 0.068 0.813 0.126 0.539 0.127 0.670
RARR 0.649 0.058 0.868 0.109 0.743 0.646 0.060 0.850 0.112 0.734 0.516 0.066 0.594 0.119 0.552 0.113 0.676
ARE 0.670 0.756 0.910 0.826 0.772 0.682 0.739 0.898 0.811 0.759 0.584 0.682 0.926 0.785 0.716 0.807 0.749
Mintaka
DRQA 0.431 0.673 - - - 0.380 0.640 - - - 0.368 0.600 - - - - -
EFEC 0.557 0.040 0.729 0.076 0.632 0.538 0.057 0.728 0.106 0.619 0.498 0.029 0.739 0.056 0.595 0.079 0.615
CCVER 0.630 0.069 0.937 0.129 0.753 0.582 0.073 0.901 0.135 0.707 0.397 0.036 0.850 0.069 0.541 0.111 0.667
RARR 0.646 0.060 0.850 0.112 0.734 0.651 0.065 0.829 0.121 0.729 0.543 0.058 0.679 0.107 0.603 0.113 0.689
ARE 0.716 0.807 0.914 0.857 0.803 0.712 0.767 0.887 0.823 0.790 0.631 0.706 0.940 0.806 0.755 0.829 0.783
StrategyQA
DRQA 0.237 0.490 - - - 0.237 0.379 - - - 0.234 0.467 - - - - -
EFEC 0.354 0.049 0.716 0.092 0.474 0.319 0.051 0.666 0.095 0.432 0.361 0.031 0.721 0.059 0.481 0.082 0.462
CCVER 0.372 0.047 0.932 0.089 0.532 0.435 0.063 0.917 0.118 0.590 0.323 0.058 0.854 0.109 0.468 0.105 0.530
RARR 0.356 0.097 0.862 0.174 0.504 0.449 0.073 0.846 0.134 0.586 0.412 0.057 0.604 0.104 0.490 0.138 0.527
ARE 0.463 0.559 0.899 0.689 0.611 0.474 0.502 0.907 0.646 0.623 0.484 0.533 0.912 0.673 0.633 0.667 0.622
ExpertQA
DRQA 0.127 0.283 - - - 0.131 0.326 - - - 0.147 0.319 - - - - -
EFEC 0.343 0.071 0.686 0.129 0.457 0.356 0.076 0.698 0.137 0.472 0.357 0.058 0.719 0.107 0.477 0.124 0.469
CCVER 0.292 0.071 0.967 0.132 0.449 0.282 0.081 0.942 0.149 0.434 0.212 0.064 0.845 0.119 0.339 0.133 0.407
RARR 0.340 0.078 0.904 0.144 0.494 0.400 0.084 0.851 0.153 0.544 0.353 0.074 0.606 0.132 0.446 0.143 0.495
ARE 0.412 0.438 0.917 0.593 0.569 0.425 0.417 0.924 0.575 0.582 0.390 0.386 0.942 0.548 0.552 0.572 0.568
5. Experiments
In this section, we outline the experimental setups, present the experimental results, and provide a thorough experimental analysis.
5.1. Evaluation Setups
This section mainly introduces the datasets used and baseline methods for comparison, and the evaluation metrics employed.
Dataset. We perform extensive experiments on three Question Answering (QA) datasets: Natural Questions (NQ) (kwiatkowskinatural, ), Mintaka (sen2022mintaka, ), and StrategyQA (geva2021strategyqa, ), as well as the AQA dataset ExpertQA. These datasets are used to evaluate the attribution ability in knowledge-intensive QA task. Following the standard dataset settings (gao2023rarr, ), we randomly select 150 samples from NQ, Mintaka and StrategyQA as test datasets. For ExpertQA, we use the entire provided test set. Details of the datasets can be found in Appendix C.1.
Baselines. The proposed ARE is compared with four post-hoc retrieval based baselines, including: EFEC [^4] https://github.com/j6mes/acl2021-factual-error-correction (thorne-vlachos-2021-evidence, ), DRQA (bohnet2022attributed, ), RARR [^5] https://github.com/anthonywchen/RARR (gao2023rarr, ) and CCVER [^6] https://github.com/jifan-chen/Fact-checking-via-Raw-Evidence (chen-etal-2024-complex, ). More details about the baselines can be found in Appendix C.2.
Metrics. We assess the attribution and editing through several metrics: $Attr_{r}$ is a molecular-level attribution metric, which measures the recall of retrieved evidences (gao2023rarr, ):
(5) $Attr_{r}(X,A)=\underset{MF_{i}\in X}{\textrm{avg}}\ \underset{E_{i}\in A}{\textrm{max}}\ \mathrm{NLI}(E_{i},MF_{i}),$
where, $MF_{i}\in X$ is a molecular clause and $E_{i}\in A$ is the evidence for $MF_{i}$, “max” selects the highest entailment score among all evidences, and “avg” calculates the average score of all evidence, $\mathrm{NLI}(E_{i},MF_{i})$ [^7] https://huggingface.co/google/t5_xxl_true_nli_mixture represents the model probability of $E_{i}$ entailing $MF_{i}$.
However, $Attr_{r}$ focuses solely on evidence recall, neglecting the precision of the recalled evidence. This can result in high scores when the quantity of evidence is sufficiently large. In light of this, we propose $Attr_{p}$ to simultaneously assess the precision and completeness of the evidences by calculating the proportion of invalid evidence among all evidences, which is defined as:
(6) $Attr_{p}=\frac{\sum_{i=1}^{m}\mathbb{I}({\textrm{NLI}_{bi}(E_{i},MF_{i})})}{m},$
where $\mathbb{I}(\mathrm{condition})$ is 1 if the condition is true, 0 for other wise; $E_{i}\in A$ and $MF_{i}\in X$. If $E_{i}$ can not entail $MF_{i}$, it was defined as invalid evidence. As for $\mathrm{NLI}_{bi}(\cdot,\cdot)$ is a binary classification result based on TRUE model. Different from the $\mathrm{NLI}(\cdot,\cdot)$, $\mathrm{NLI}_{bi}(\cdot,\cdot)$ will return true only when the evidence supports the whole sentence, which makes $Attr_{p}$ more stricter than $Attr_{r}$. The proof of $Attr_{p}$ can be used to evaluate the precision is available in Appendix E.
Preservation (gao2023rarr, ) generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$:
(7) $Pres_{(X,X^{\prime})}=\textrm{max}(1-\frac{\textrm{Lev}(X,X^{\prime})}{\textrm{length}(X)},0),$
where $Pres_{(X,X^{\prime})}$ equals 1 when $X$ is identical to $X^{\prime}$, indicating no changes. A value of 0 means $X$ and $X^{\prime}$ share no common words, reflecting complete divergence.
To better compare with different baselines, $F1_{RP}$ (gao2023rarr, ; chen2023purr, ) and $F1_{PP}$ is proposed. $F1_{RP}$ and $F1_{PP}$ are calculated by the following equations:
[…]
More details of the metrics can be found in Appendix C.3.
Table 2. The ablation experiment results on NQ, StrategyQA, and ExpertQA using GPT-3.5. “w/o” represents removal for the mentioned module. The ablation study on Mintaka is shown in Appendix C.2. We mark the better results in bolded.
Methods NQ StragegyQA ExpertQA
[…]
w/o edit 0.648 0.715 - - - 0.442 0.53 - - - 0.394 0.417 - - -
w/o atomic 0.451 0.619 0.9 0.734 0.608 0.225 0.331 0.918 0.487 0.362 0.168 0.308 0.943 0.464 0.285
w/o molecular 0.713 0.785 0.585 0.670 0.643 0.514 0.609 0.598 0.603 0.553 0.425 0.449 0.561 0.499 0.484
w/o re-retrieval 0.609 0.677 0.911 0.777 0.730 0.437 0.533 0.927 0.677 0.594 0.388 0.421 0.937 0.581 0.548
ARE 0.69 0.737 0.91 0.814 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
5.2. Experimental Results and Analysis
The experimental results are displayed in Table 1, and the experimental analyses are listed as follows:
(1) ARE demonstrates significant improvements in both molecular-level attribution evaluation metrics and intent preservation across four datasets for all LLMs. ARE achieves improvements of 68%, 71.6%, 52.9%, and 42.9% over the $Ave-\mathrm{F1_{PP}}$ metric across all datasets. Additionally, ARE also achieves the best performance in the $Ave-\mathrm{F1_{RP}}$ metric, with improvements of 7.3%, 9.4%, 9.2%, and 7.3% on GPT-3.5, Llama3-70B, and Llama2-70B [^8] We use GPT-3.5-turbo-1106 for reproducibility. The Llama series models can be downloaded from https://huggingface.co/meta-llama/Llama-2-70b-chat and https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct., respectively, demonstrating the superior effectiveness and generalization capabilities of our proposed ARE.
(2) For attribution ability, ARE outperforms other baselines on $Attr_{r}$ and $Attr_{p}$. Especially in $Attr_{p}$, all existing methods are far below ARE, this is because $Attr_{p}$ is a strict metric that evaluates the completeness of evidence. As for Pres, although the CCVER gets a higher $Pres$ in GPT-3.5 and Llama3-70B, it performed poorly on Llama-2-70B. In contrast, ARE achieved the most stable intent preservation across all LLMs, even with the less effective Llama2-70B model, which struggles with intent preservation, particularly when using the RARR method. These results demonstrate that the atomic fact-base ARE can maximize the attribution score while preserving the original intent.
(3) The sub-question generation methods perform similarly on NQ, Mintaka, and StrategyQA, but CCVER significantly underperforms compared to RARR on ExpertQA. This highlights the limitations of sub-question generation, which require carefully designed prompts. Conversely, ARE shows greater robustness and adaptability across different datasets. Further experimental results are available in Appendix D.
Figure 4. The performance of the fact decomposition LLM at the molecular level and the atomic level for different number of iterations.
5.3. Ablation Study
In order to investigate the impact of key modules on experimental performance, we conduct a series of ablation experiments, and the corresponding results are presented in Table 2. Specifically, “w/o edit” means removing the editing module; “w/o atomic” represent removing atomic-level facts; “w/o molecular” represents removing molecular-level clause; “w/o re-retrieval” means removing the status of “irrelevant” in evidence verifier.
(1) In the “$w/o$ edit” setting, attribution scores decline significantly. For example, $Attr_{r}$ decreased by 4.2%, while $Attr_{p}$ dropped by 2.2% in NQ. These results underscore the importance of atomic fact editing, as modifying hallucinated content is crucial for further improving attribution accuracy.
(2) The results of the “w/o atomic” setting show the atomic facts play an important role in attribution scores. Specifically, ARE obtains 23.9% in $Attr_{r}$ and 11.8% in $Attr_{p}$ improvements on the NQ dataset. This phenomenon indicate that atomic facts have higher certainty and lower entropy, making it possible to retrieve relevant evidence more effectively.
(3) Although the “w/o molecular” setting performs better on $Attr_{r}$ and $Attr_{p}$, it shows poorer performance on the $pres$ metric. This may be attributed to the removal of molecular facts, which causes atomic facts to lose their correspondence with them, preventing accurate backtracking to their original positions for editing. Consequently, this significantly increases the risk of altering the original intent.
(4) The performance of “w/o re-retrieval” shows that the necessity of the “irrelevant” status in evidence verifier module. For example, the $Attr_{r}$ decreased by 8.1% and 6% in $Attr_{p}$ in the NQ dataset. This phenomenon demonstrates that expanding facts when they have irrelevant evidence and then re-retrieving evidence can effectively enhance the attribution capability.
Table 3. Comparison of the proposed atomic fact-based retrieval with previous sub-question-based retrieval performance.
Methods NQ StrategyQA ExpertQA
$Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$
DRQA 0.424 0.647 0.237 0.49 0.127 0.283
CCVER 0.602 0.044 0.359 0.044 0.296 0.073
RARR 0.593 0.043 0.302 0.096 0.305 0.077
ARE 0.648 0.715 0.442 0.550 0.404 0.437
[…]
5.4. The Impact of Fact Decomposition LLM with Different Iterations
To find the best performance of molecular-to-atomic fact decomposition LLM and explore how iterations affect fact decomposition performance, we perform a comparative experiment, and the results are illustrated in Figure 4. Specificlly, we assess its performance at both the molecular and atomic levels, focusing on two key aspects: consistency and correctness. Consistency $n_{c}$ measures how closely the number of decomposed sentences aligns with that of gold sentences. Correctness $d_{correct}$ evaluates whether the decomposed sentences preserve the original sentence meaning.
We test on a well-constructed evaluation dataset mentioned in 4.1 to select the decomposition model used in ARE. As shown in Figure 4, the performance of $n_{c}$ at the molecular-level initially increase and then decreasing as the number of epochs increase. At the atomic-level, $n_{c}$ shows an overall upward trend, stabilizing at epoch 4. For $d_{correct}$, performance at molecular-level also increase and then decrease as the number of epochs increase, peaking at epoch 3 with 0.972. However, at the atomic-level, it first decreases and then increases.
In terms of consistency, the performance at the 3rd and 4th epochs is similar. However, for correctness, the score of the atomic-level at epoch 3 is the highest, and meanwhile the score of the molecular-level at epoch 3 and epoch 4 has little difference. Therefore, we select the LLM trained with 3 epochs. Details of evaluation process can be found in Appendix D.5.
Table 4. Comparing the revised results with different systems. Red indicates hallucinated content that conflicts from the evidence or fact. Blue marks the content that, after revision, has been corrected to align with the evidence. Orange indicates it is still not correct after the revision.
Original long-form answer
”Daddy’s Little Girl” is a song by American country music artist Kellie Pickler. The song was released in 2006 as the second single from her debut album ”Small Town Girl.” Gold Answer: “Al Martino”
Revised by EFEC
”The Daddy’s Little Girl is a book written and recorded by a Canadian music producer, and author.”       Pres: 0.322
Revised by RARR
”Daddy’s Little Girl” is a classic song typically played at white weddings while a bride dances with her father. The song’s lyrics and music were first written by Robert Harrison Burke and Horace Gerlach in 1949.       Pres: 0.245
Revised by ARE
”Daddy’s Little Girl” is a song by American traditional pop and jazz singer Al Martino. The song was released in 2006 as the second single from her debut album ”Small Town Girl.” Pres: 0.83
5.5. Hypothesis Proof Experiment
As illustrated in Table 3, ARE shows the improvements of 4.6% on NQ, 8.3% on StrategyQA, and 9.9% on ExpertQA based on the metric of $Attr_{r}$. This supports the hypothesis that atomic facts are easier to retrieve as relevant evidence. Furthermore, DRQA, using the entire answer and question for retrieval, struggles in cross-document retrieval scenarios. Methods involving generated sub-questions fail to obtain more evidence due to the poor quality of generated questions.
CCVER and RARR both generate sub-questions and then use these sub-questions for retrieval. Their performance is similar on NQ and ExpertQA, but on StrategyQA, CCVER outperforms RARR. This is because CCVER generates “Yes or No” type questions, which match the type of the StrategyQA dataset, allowing for better evidence retrieval. However, both perform worse than ARE, which is based on atomic fact retrieval.
5.6. Visualization results of attribution scores and preservation
To provide a more intuitive comparison between existing methods and our proposed ARE, regarding the simultaneous achievement of maximum attribution scores and preservation of original intent, we conducted the visualization experiments illustrated in Figure 5.
As shown in Figure 5, ARE is positioned on the right side of the dashed line, which indicates that ARE has achieved effective editing. The contour lines indicate that the proposed ARE achieves the best performance in both the F1RP (F1RP both considers Attrr and $Pres$) and F1PP (F1RP simultaneously considers Attrp and $Pres$) metrics after editing. These phenomena demonstrates the robustness of the proposed ARE in maximizing the attribution scores and preserving the original intent. The completed results can be found in Appendix D.4.
5.7. Case Study
When comparing the revised results of EFEC, RARR, and ARE in Table 4, clearly highlights its advantages, as well as the shortcomings of RARR and EFEC. ARE performs minimal edits by focusing on atomic facts, successfully preserving the original intent while improving factuality and attribution. In contrast, RARR often makes substantial modifications, which can make it challenging to preserve the original intent of the content. EFEC performs even less satisfactorily. It makes significant changes to the original content, omitting entire sentences. Such drastic revisions can result in a final product that significantly deviates from the original content and intent.
6. Conclusion
This paper proposes a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for post-hoc retrieval-based AQA tasks, which contains a Molecular-to-Atomic decomposition stage and an Atomic-to-Molecular backtracking process. Specifically, ARE employs an instruction-tuned fact decomposition LLM to decompose the long-form answers into multiple molecular clauses and atomic facts. This LLM is fine-tuned in a well-constructed molecular-to-atomic fact decomposition dataset. Subsequently, ARE leverages an LLM-based verifier to validate the relationships between the searched evidences and their corresponding atomic facts. Based on the verifier’s assessment, ARE determines whether the facts require further expansion for re-retrieval or editing. Furthermore, ARE proposes a more comprehensive evaluation metric $Attr_{p}$, which not only accurately measures the precision of retrieved evidence, but also emphasizes the completeness of the evidence. The effectiveness of this framework is demonstrated across four datasets using four prominent LLMs, complemented by an extensive ablation study and LLM evaluation.
Appendix
Appendix A Prompt for LLMs
A.1. Prompt for Long-form answer generation
Prompt for Long-form answer generation you need think step by step and answer my question. Example 1: 1. Question: Which actor was the star of Titanic and was born in Los Angeles, California? 2. Explanation: Titanic is a film directed by James Cameron. Leonardo DiCaprio was played the character Jack Dawson in Titanic. Leonardo DiCaprio was born in Los Angeles,California. 3. Answer: Leonardo DiCaprio. Example 2: 1. Question: How many teams has Matthew Stafford played for? 2. Explanation: Matthew Stafford is an American football quarterback. Matthew Stafford played for the Lions from 2009 to 2020. Matthew Stafford played for Los Angeles Rams in 2021. 3. Answer: 2. Example 3: 1. Question: Is Jake Gyllenhaal older than Maggie Gyllenhaal? 2. Explanation: Jake Gyllenhaal is an American actor. Jake Gyllenhaal was born in December 19, 1980, and Maggie Gyllenhaal is an American actress. Gyllenhaal was born in November 16, 1977. 3. Answer: No. Example 4: 1. Question: How many sacks does Clay Matthews have in his career? 2. Explanation: Clay Matthews has been pro linebacker for 11 seasons. Ten of the seasons Clay Matthews played for Green Bay Packers,while Clay Matthews last season was with Los Angeles Rams, dedicated to Clay Matthews hometown. Clay Matthews III had 383 solo tackles, 136 assists, and 91.5 sacks. Clay Matthews father, Clay Matthews Jr., also played in NFL and had 69.5 sacks in his career. 3. Answer: 91.5 (Clay Matthews III) and 69.5 (Clay Matthews Jr.). please answer the Question: question and give the Explanation and a answer start with the ”Explanation” and ”Answer”.
A.2. An example instruction data for instruction-tuning
An example data for instruction-tuning {“instruction”: “Please extract the clauses and the corresponding atomic facts from the following text, when the pronoun appears, output the real name of the person or thing it refers to.”, input”: “Lawrence Wackett’s occupation is an aerospace engineer. He was involved in World War I and received the James Cook Medal on January 1, 1978. Lawrence Wackett’s gender is male.”, output”: [{“Lawrence Wackett’s occupation is an aerospace engineer.”: [“Lawrence Wackett’s occupation is an aerospace engineer.”]}, {’He was involved in World War I and received the James Cook Medal on January 1, 1978.’: [’Lawrence Wackett was involved in World War I.’, ’Lawrence Wackett received the James Cook Medal on January 1, 1978.’]}, {“Lawrence Wackett’s gender is male.”: [“Lawrence Wackett’s gender is male.”] }] }
A.3. Prompt for evidence verifing and fact expansion
Prompt for evidence verifing I will check some things you said. Here are some examples for you to learn this process: Example 1: 1. You said: Your nose switches back and forth between nostrils. When you sleep, you switch about every 45 minutes. This is to prevent a buildup of mucus. It’s called the nasal cycle. 2. I checked: When you sleep, you switch about every 45 minutes. 3. I found this article: Although we don’t usually notice it, during the nasal cycle one nostril becomes congested and thus contributes less to airflow, while the other becomes decongested. On average, the congestion pattern switches about every 2 hours, according to a small 2016 study published in the journal PLOS One. 4. Reasoning: The article said the nose’s switching time is about every 2 hours, and you said the nose’s switching time is about every 45 minutes. 5. Therefore: This disagrees with what you said. Example 2: 1. You said: The Little House books were written by Laura Ingalls Wilder. The books were published by HarperCollins. 2. I checked: The Little House books were published by HarperCollins. 3. I found this article: These are the books that started it all – the stories that captured the hearts and imaginations of children and young adults worldwide. Written by Laura Ingalls Wilder and published by HarperCollins, these beloved books remain a favorite to this day. 4. Reasoning: The article said the Little House books were published by HarperCollins and you said the books were published by HarperCollins. 5. Therefore: This agrees with what you said. Example 3: 1. You said: Real Chance of Love was an American reality TV show. Season 2 of the show was won by Cali, who chose to be with Chance. 2. I checked: Season 2 of the show was won by Cali. 3. I found this article: Real Chance of Love 2: Back in the Saddle is the second season of the VH1 reality television dating series Real Chance of Love. Ahmad Givens (Real) and Kamal Givens (Chance), former contestants on I Love New York are the central figures. 4. Reasoning: The article doesn’t answer the question and you said that Cali won season 2 of Real Chance of Love. 5. Therefore: This is irrelevant to what you said. Example 4: 1. You said: Tiger Woods is the only player who has won the most green jackets. He has won four times. The Green Jacket is one of the most coveted prizes in all of golf. 2. I checked: The Green Jacket is one of the most coveted prizes in all of golf. 3. I found this article: The green jacket is a classic, three-button, single-breasted and single-vent, featuring the Augusta National Golf Club logo on the left chest pocket. The logo also appears on the brass buttons. 4. Reasoning: The article said the Green Jacket is a classic three-button single-breasted and single-vent and you said the Green Jacket is one of the most coveted prizes in all of golf. 5. Therefore: This is irrelevant to what you said. Now, please follow above examples pattern to inference new one (don’t output the examples), and outputs the result start with ’4. Reasoning: ’ and ’5. Therefore: ’. If from the article supports about ’you said’, you should generate ”agrees”, if unsupports, you shuld generate ”disagrees” and if the article is not fully support ’you said’, you should generate ”irrelevant” in ’Therefore’. 1. You said: {claim 2. I checked: {query} 3. I found this article: {evidence}
Prompt for fact expansion Rewrite the fact: {fact} into two short atomic phrases based on Wikipedia. Ensure they contain more factual information and can be easily supported by search engines.
A.4. Prompt for editing
Prompt for atomic editing you need use some evidences to check ’You said’ if there are some difference and generate in ”My fix”. Here are some examples to learn: Example 1: 1. You said: 11 Diagonal Street was actually completed way back in 1978. 2. I found these evidences: 11 Diagonal Street completion date ”1983” . 11 Diagonal Street building end date ”1983” . 11 Diagonal Street building start date ”1978”. 3. This suggests 11 Diagonal Street was completed in 1978 is wrong. 4. My fix: 11 Diagonal Street was actually completed way back in 1983. Example 2: 1. You said: The length of Mississippi River is 2340 km. 2. I found these evidences: Missouri River length +3726. Mississippi River discharge +12743. Mississippi River tributary Yazoo. 3. The question is What is the length of the Mississippi River? This suggests 2340 km of Mississippi River in your statement is wrong. 4. My fix: The length of Mississippi River is 3726 km. Example 3: 1. You said: Your nose switches back and forth between nostrils. When you sleep, you switch about every 45 minutes. This is to prevent a buildup of mucus. It’s called the nasal cycle. 2. I found this evidences: Although we don’t usually notice it, during the nasal cycle one nostril becomes congested and thus contributes less to airflow, while the other becomes decongested. On average, the congestion pattern switches about every 2 hours, according to a small 2016 study published in the journal PLOS One. 3. This suggests 45 minutes switch time in your statement is wrong. 4. My fix: Your nose switches back and forth between nostrils. When you sleep, you switch about every 2 hours. This is to prevent a buildup of mucus. It’s called the nasal cycle. Example 4: 1. You said: In the battles of Lexington and Concord, the British side was led by General Thomas Hall. 2. I found this evidences: Interesting Facts about the Battles of Lexington and Concord. The British were led by Lieutenant Colonel Francis Smith. There were 700 British regulars. 3. This suggests General Thomas Hall in your statement is wrong. 4. My fix: In the battles of Lexington and Concord, the British side was led by Lieutenant Colonel Francis Smith. Now, I will give you a new instance, please follow the above example to fix new one and start with ”My fix:”, do not generate irrelevant information. 1. You said: {claim} 2. I found these evidences: {evidence} 3. This suggests
Appendix B Theoretical Analysis of the Fact Expansion
When the retrieved evidence is invalid, we consider the possible reasons to be: The informativeness of the query contained may too lower. Increasing the informativeness of the query is an avenue that can be tried, and and the process will be justified in terms of informativeness. The informativeness of an atomic fact can be defined as:
$I(AF_{ij})$ $=-\log P(\boldsymbol{AF}_{ij})$
$=P(t_{1},\dots,t_{i},\dots,t_{n})$
$=P(t_{1})\cdot\dots\cdot P(t_{n}\mid t_{1},t_{2},\dots,t_{n-1}),$
where $t_{i}$ is a token in $AF_{ij}$. We assume that the set of facts that can be extended from $AF_{ij}$ is $S_{ex}=\{ef_{1},...,ef_{j},...,ef_{n}\}$, then the information content of each $ef_{j}$ is $I(ef_{j})$. Finally, for an invalid query, we expand it to $AF_{ij}^{ex}$, then its final information content is
$I(AF_{ij}^{ex})$ $=-\log P(AF_{ij})-\log P(ef_{i})$
$=I(AF_{if})+I(ef_{j})>I(AF_{if}).$
Table 5. The experimental results of aggregating evidence retrieved from sub-questions into the corresponding molecular clauses.
GPT-3.5 NQ StrategyQA ExpertQA
[…]
RARR 0.649 0.058 0.868 0.109 0.743 0.356 0.097 0.862 0.174 0.504 0.34 0.078 0.904 0.144 0.494
RARR w aggregate evidnece 0.499 0.646 0.942 0.766 0.653 0.336 0.53 0.948 0.680 0.496 0.276 0.407 0.959 0.571 0.428
ARE 0.67 0.756 0.91 0.826 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
[…]
Appendix C The details of the Experimental Settings
C.1. Datasets
Natural Questions is a question-answering dataset developed by the Google search engine and is widely used to evaluate machine reading comprehension, information retrieval, and other tasks. Its questions are complex and often ambiguous, requiring the identification of relevant answers from long documents, making it a knowledge-intensive task.
Mintaka is a complex multilingual dataset featuring superlative, cross, and multi-hop question types. It is based on knowledge graphs (KG), which are also knowledge-intensive datasets widely used to evaluate end-to-end deep learning question-answering models.
StrategyQA is a dataset focused on open-domain problems, which involve inference steps in the questions and require the model to provide the reasoning process to arrive at the correct answer. This places a high demand on the model’s attribution ability, as its answers typically reference different paragraphs from multiple documents.
ExpertQA is a dataset specifically designed to evaluate attribution-based question answering models. It encompasses seven question types across 32 domains, and answering these questions requires specialized knowledge, posing a greater challenge for attribution systems.
C.2. Baselines
  • DRQA [8] is a straightforward attribution model that concatenates the question and answer into a single query before performing retrieval. The purpose of this baseline is to demonstrate how much data in the dataset does not require decomposition prior to retrieval.
  • RARR [9] is the first method to implement a retrieval-then-revise paradigm in the AQA task, with all its components based on API-driven LLMs. It first prompts the LLM to generate a long-form answer as a claim, then generates a series of sub-questions based on this claim, using these sub-questions as queries to retrieve evidence from a search engine. RARR has also shown that using Bing Search and Google Search as evidence sources yields similar performance.
  • EFEC [41] is an editor built on the T5 model that modifies a claim by incorporating evidence. It is trained using both weak and full supervision methods. To demonstrate its optimal performance, we reproduced it using full supervision. EFEC employs a deep neural network-based retriever for evidence retrieval. For a fairer comparison, we adopted the same settings as RARR, generating sub-questions for retrieval and using the retrieved evidence to edit the answer.
  • CCVER [42] is a model focused on fact verification, consisting of a two-stage retrieval process. In the first stage, it uses a few-shot approach to prompt LLMs to generate Yes or No-type sub-questions for a given claim, which are then used as queries to retrieve relevant evidence. In the second stage, it combines BM25 and LLMs to summarize the evidence. Finally, DeBERTa-large is employed to evaluate the claim based on the evidence. We selected CCVER to explore the impact of generating different types of questions on attribution performance, utilizing only its question generation prompt in conjunction with RARR’s editing method.
Since the models reported in the RARR paper (GPT-3 and PaLM) are no longer available, we used other LLMs to reproduce RARR.
Table 6. Ablation experimental results of GPT3.5 on Mintaka
Methods Mintaka
[…]
w/o edit 0.672 0.768 - - -
w/o atomic 0.477 0.651 0.928 0.765 0.630
w/o molecular 0.725 0.832 0.548 0.661 0.624
w/o re-retrieval 0.680 0.767 0.936 0.843 0.788
ARE 0.716 0.807 0.914 0.857 0.803
C.3. Metric
NLI model. We use an NLI model based on T5-11B, which is trained on six datasets: MNLI, SNLI, FEVER, PAWS, SciTail, and VitaminC  [43, 44, 45, 46, 47, 48]. The input format is: “premise: {evidence} hypothesis: {sentences}”. As suggested by [9], we use the probability of producing “1” as the entailment score in the metric $Attr_{r}$ and use the binary result “1” (entailed) or “0” (not entailed) for $Attr_{p}$.
Levenshtein distance. Preservation generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$, which assesses the difference between two strings. Levenshtein distance is defined as the minimum number of single-character edits (insertions, deletions, or substitutions) required to transform one string into another. This metric quantifies the dissimilarity between two sequences and is commonly used in applications such as spell-checking and natural language processing.
C.4. Other Settings in Experiments.
The Complete Heuristics Rules
The complete heuristics rules used in instruction-tuning dataset construction Triples containing the following will be filtered: “ID”, “.svg”, “.jpg”, “.png”, “.JPG”, “.JPEG”, “http://”, “https://”, “Category:”, “Wikipedia:”, “code”,“UMLS CUI”, “Wikimedia”, “.map”,“ISO”,“code”, “UNESCO”,“html” , “IPTC” ,“ISNI” , “ISSN”.
Molecular-to-Atomic fact decomposition LLM. To instruction-tune the Molecular-to-Atomic fact decomposition LLM, we used Llama-Factory [^9] https://github.com/hiyouga/LLaMA-Factory [49] for fine-tuning. The hyperparameters used in the training are: The number of epochs is 3, learning rate is 1e-4, batch_size is 1, warmup_ratio is 0.1, cutoff_len is 2048, per_device _train_batch_size is 1 and gradient_accumulation_steps is 2, lr _scheduler_type is cosine. The entire training and inference are conducted on 2 * A800 80G GPUs.
LLMs. To efficiently infer the LLaMA series of large language models, we utilized VLLM [^10] https://github.com/vllm-project/vllm for inference on two A800 GPUs. To ensure reproducibility and stability, all temperature settings are set to 0, except for the Llama3 series, for which we adjusted the temperature to the lowest possible value of 0.1, as it cannot be set to 0.
Table 7. Ablation experiment results based on GPT-4o-mini
[…]
GPT-4o-mini NQ StragegyQA
w/o edit 0.587 0.706 - - - 0.364 0.467 - - -
w/o atomc 0.328 0.541 0.884 0.671 0.479 0.189 0.237 0.883 0.374 0.280
w/o molecular 0.653 0.72 0.544 0.620 0.594 0.429 0.514 0.563 0.537 0.487
w/o re-retrieval 0.602 0.691 0.924 0.791 0.729 0.402 0.462 0.914 0.614 0.558
ARE 0.623 0.722 0.902 0.802 0.737 0.413 0.469 0.904 0.618 0.567
Mintaka ExpertQA
w/o edit 0.634 0.766 - - - 0.377 0.401 - - -
w/o atomc 0.349 0.566 0.870 0.686 0.499 0.195 0.290 0.921 0.441 0.321
w/o molecular 0.648 0.76 0.446 0.562 0.528 0.391 0.415 0.549 0.473 0.457
w/o re-retrieval 0.619 0.707 0.871 0.780 0.723 0.369 0.390 0.939 0.551 0.529
ARE 0.665 0.755 0.881 0.813 0.738 0.387 0.414 0.914 0.570 0.545
Appendix D Other experimental results
D.1. The Ablation Study on Mintaka with GPT-3.5
As shown in Table 6, similar to the other datasets, each module in ARE enhances attribution, particularly the editing module, which increases $Attr_{r}$ by 0.044 and $Attr_{p}$ by 0.039 on the Mintaka dataset.
D.2. Experimental results on GPT-4o-mini.
We also conducted experiments on the latest GPT series model, GPT-4o-mini (To ensure reproducibility, we used GPT-4o-mini-2024-07-18.). The experimental results can be found in Table 8.
We also conducted ablation experiments on GPT-4o-mini, and the results can be found in Table 7. The experimental results show that ARE also achieved the best performance on GPT-4o-mini, maximizing the attribution score while preserving the original intent.
Table 8. Experimental results using GPT-4o-mini on four datasets.
GPT-4o-mini
[…]
NQ
DRQA 0.265 0.576 - - -
EFEC 0.452 0.053 0.688 0.098 0.546
CCVER 0.563 0.051 0.931 0.097 0.702
RARR 0.493 0.08 0.904 0.147 0.638
ARE 0.623 0.722 0.902 0.802 0.737
Mintaka
DRQA 0.251 0.544 - - -
EFEC 0.497 0.049 0.698 0.092 0.581
CCVER 0.634 0.045 0.933 0.086 0.755
RARR 0.508 0.058 0.899 0.109 0.649
ARE 0.634 0.755 0.881 0.813 0.737
StrategyQA
DRQA 0.08 0.222 - - -
EFEC 0.33 0.071 0.638 0.128 0.435
CCVER 0.341 0.056 0.961 0.106 0.503
RARR 0.292 0.079 0.916 0.145 0.443
ARE 0.413 0.469 0.904 0.618 0.567
ExpertQA
DRQA 0.122 0.308 - - -
EFEC 0.297 0.075 0.635 0.134 0.405
CCVER 0.31 0.062 0.98 0.117 0.471
RARR 0.269 0.064 0.944 0.120 0.419
ARE 0.387 0.414 0.914 0.570 0.544
D.3. Experimental results of aggregating evidence from sub-questions.
The sub-question generation method (such as RARR) lacks a clear correspondence between questions and clauses, failing to provide complete evidence for molecular clauses. Therefore, we employ our Molecular-to-Atomic fact decomposition LLM to first decompose the long-form answer. Next, based on the molecular clauses, we utilize the sub-question generation method from RARR to prompt LLMs to generate sub-questions for retrieval. Finally, the evidence retrieved through the sub-questions is aggregated according to its relationship with the molecular clauses, serving as supporting evidence for these clauses.
As shown in Table 5, when evidence was aggregated for molecular clauses, the $Attr_{p}$ score improved significantly, increasing by 0.588 on NQ, 0.433 on StrategyQA, and 0.329 on ExpertQA. However, the $Attr_{r}$ metric is lower than that of RARR. This is because considering each molecular clause individually can lead to a loss of context, especially crucial subjects, resulting in unclear sub-questions that fail to retrieve key information.
D.4. Visualization results of attribution scores and Pres.
The visualization results for Mintaka and ExpertQA are presented in Figure 6. These figures illustrate the distribution of performance metrics, showing how different methods, including ARE, perform in terms of attribution and preservation of intent across these two datasets.
D.5. The Evaluation process of Fact Decomposition LLM
we define two metrics: $n_{c}$ for consistency and $d_{correct}$ for correctness.
For consistency, it checks whether the number of decomposed sentences $s_{n}$ matches the number of gold sentences $g_{n}$. The decomposition is considered consistent when the number of decomposed sentences and gold sentences are equal. Formally, $n_{c}$ is defined as:
$n_{c}=\mathbb{I}(\frac{s_{n}}{g_{n}}=1),$
where $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if the condition is satisfied (i.e., the ratio equals 1) and 0 otherwise. This metric ensures that the decomposition does not produce too few or too many sentences relative to the gold reference.
For correctness, it is evaluated whether the meaning of the decomposed sentences remains consistent with the gold sentences.
Let $d_{s}$ represent a decomposed sentence and $g_{s}$ a gold sentence. The correctness assessed using a binary NLI result with TRUE model, to check if $d_{s}$ entails $g_{s}$ and vice versa, it can be defined as:
$d_{correct}=\frac{NLI_{bi}(d_{s},g_{s})}{s_{n}}.$
Here, $NLI_{bi}(d_{s},g_{s})$ measures the entailment between decomposed and gold sentences. The score is normalized by the total number of decomposed sentences $s_{n}$ , ensuring that we account for the number of decomposed sentences when calculating correctness. This metric ensures that each decomposed sentence accurately reflects its corresponding gold sentence in terms of meaning.
Appendix E The Proof of $Attr_{p}$ can be used to Evaluate the Precision
The Precision can be defined as:
Precision=True Positive (TP)True Positive (TP)+False Positive ($\text{Precision}=\frac{\text{True Positive (TP)}}{\text{True Positive (TP)}+\text{False Positive (FP)}}.$
In this paper, “True Positive (TP)” is $E_{i}$ can correctly entail at least one molecular clause $MF_{i}$, “False Positive (FP)” indicates $E_{i}$ do not entail any of the molecular clauses $MF_{i}$. According to the equation 6, $\text{NLI}_{bi}(E_{i},MF_{i})=1$ if $E_{i}$ entails $MF_{j}$, and otherwise. for each $E_{i}$, the number of FP can be calculated:
$\text{FP}=\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=0),$
For TP, it can be calculated as:
$\text{TP}=\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=1),$
The attribution report collection $\{E_{1},...,E_{m}\}$ has $m$ evidence, each of them considered as a predicted positive instance. Therefore, the total number of predicted positives is:
Total Predicted $\text{Total Predicted Positives}=\text{TP}+\text{FP}=m.$
The number of TP is: $\text{TP}=m-\text{FP}$. Applying the precision formula:
Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}}$
$=\frac{\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=1)}{m},$
due $\text{NLI}_{bi}$ will return the binary result, it can be further simplified to:
Precision $=\frac{\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i}))}{m}$
$=Attr_{p}.$

### v2

@@
 Atomic Fact Decomposition Helps Attributed Question Answering
-Abstract.
+Abstract
 Attributed Question Answering (AQA) aims to provide both a trustworthy answer and a reliable attribution report for a given question.
@@
 Extensive evaluations demonstrate the superior performance of our proposed method over the state-of-the-arts on several datasets, with an additionally proposed new metric $Attr_{p}$ for evaluating the precision of evidence attribution.
-Attributed Question Answer, Information Retrieval, Large Language Models
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; April 28 - May 2; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 184ccs: Information systems Information retrievalccs: Computing methodologies Natural language generation
-1. Introduction
-Large Language Models (LLMs), pre-trained on large-scale text corpora (10387715, ), have demonstrated remarkable capabilities in natural language understanding and generation tasks (zhuang2023toolqa, ; wang2024large, ).
-However, despite their impressive performance, they often face significant challenges in real-world applications due to issues like lack of interpretability and indecisiveness (huang2023survey, ; peskoff-stewart-2023-credible, ).
-These limitations undermine the reliability and trustworthiness of LLMs, particularly in contexts where transparency and accuracy are essential.
-To address these issues, Attributed Question Answering (AQA) (li2023survey, ) has emerged as a solution, focusing on linking their generated answers to specific sources of evidence.
-Figure 1.
-The motivation of previous methods and our proposed ARE.
+Index Terms:
+Attributed Question Answer, Information Retrieval, Large Language Models.
+I Introduction
+Large Language Models (LLMs), pre-trained on large-scale text corpora [1], have demonstrated remarkable capabilities in natural language understanding and generation tasks [2, 3].
+However, their application in real-world scenarios is hindered by a lack of transparency, limited interpretability, and occasional factual inaccuracies, [4, 5].These limitations compromise the reliability of LLM outputs, especially in settings where factual correctness and evidence traceability are critical.
+To address these challenges, Attributed Question Answering (AQA) [6] has emerged as a promising paradigm.
+AQA aims to enhance answer trustworthiness by explicitly linking generated responses to verifiable sources of evidence.
+As illustrated in Figure 1, when asked “Who has the most Super Bowl rings?”, an LLM initially responds: “The player with the most Super Bowl rings is Tom Brady.
+Tom Brady is an American football quarterback who has won six Super Bowl championships.” Upon retrieving external evidence, however, the system identifies a factual error — Tom Brady has won seven, not six, championships.
+This triggers a post-hoc correction process, resulting in a revised and properly attributed answer, along with the supporting evidence.
+Depending on the timing of retrieval, previous research can be divided into two categories: Retrieval-then-Read (RTR)  [7], and Post-hoc Retrieval [8, 9, 10, 11].
+RTR delivers relevant answers by retrieving documents based on the query, providing detailed context that can enhance the richness of the response.
+However, LLMs often link content to irrelevant or incorrect sources and lack an effective mechanism for subsequent correction.
+In contrast, post-hoc retrieval typically retrieves specific external information for initially generated long-form answers, effectively reducing linking errors and enabling targeted revisions, which enhances flexibility and robustness [12, 13].
+Despite the flexibility post-hoc retrieval offers in verifying each fact and correcting it in long-form answers, it introduces several significant challenges.
+Firstly, existing methods [10, 13] prompt LLMs to directly generate a series of sub-questions for retrieval based on the long-form answer (As shown in Figure 1 (I)).
+However, the complex logical structures and vast amounts of information typically found in long-form answers make it difficult for these methods to capture all necessary facts comprehensively.
+As a result, the generated sub-questions often fail to align with the core content of the answer, leading to irrelevant questions and ultimately hindering the retrieval of accurate supporting evidence.
+Moreover, existing methods [9, 14] that directly edit long-form answers holistically often face difficulty in precisely locating specific content for revision, making it challenging to preserve the original intent while ensuring consistency between the answer and the attribution.
+This limitation arises from either insufficient editing, where critical details remain inadequately revised, or excessive editing, which distorts the original intent or introduces unnecessary changes, which disrupts the coherence of the answer.
+Thus, effectively balancing the challenges of insufficient and excessive editing has become a critical issue.
+Figure 1: The motivation of previous methods and our proposed ARE.
 Previous methods directly use LLMs to generate sub-questions for the long-form answer and edit the whole answer.
 Our proposed ARE leverages instruction-tuned LLMs to achieve molecular-to-atomic stage, and backtrack the edited atomic facts to achieve atomic-to-molecular stage.
-Depending on the timing of retrieval, previous research can be divided into two categories: Retrieval-then-Read (RTR)  (gao2023enabling, ), and Post-hoc Retrieval (bohnet2022attributed, ; gao2023rarr, ; chen2023purr, ).
-RTR delivers relevant answers by retrieving documents based on the query, providing detailed context that can enhance the richness of the response.
-However, LLMs often link content to irrelevant or incorrect sources and lack an effective mechanism for subsequent correction.
-In contrast, post-hoc retrieval typically retrieves specific external information for initially generated long-form answers, effectively reducing linking errors and enabling targeted revisions, which enhances flexibility and robustness.
-(dhuliawala2023chainofverification, ; kim2024re, ).
-Despite the flexibility post-hoc retrieval offers in verifying each fact and correcting it in long-form answers, it introduces several significant challenges.
-Firstly, existing methods (chen2023purr, ; kim2024re, ) prompt LLMs to directly generate a series of sub-questions for retrieval based on the long-form answer (As shown in Figure 1 (I)).
-However, the complex logical structures and vast amounts of information typically found in long-form answers make it difficult for these methods to comprehensively capture all necessary facts.
-As a result, the generated sub-questions often fail to align with the core content of the answer, leading to irrelevant questions and ultimately hindering the retrieval of accurate supporting evidence.
-tribute methods always overlooking the broader context, which can lead to inaccurate or incomplete attribution.
-Moreover, existing methods (gao2023rarr, ; wu2024updating, ) that directly edit long-form answers holistically often face difficulty in precisely locating specific content for revision, making it challenging to preserve the original intent while ensuring consistency between the answer and the attribution.
-This limitation arises from either insufficient editing, where critical details remain inadequately revised, or excessive editing, which distorts the original intent or introduces unnecessary changes, which disrupts the coherence of the answer.
-Thus, effectively balancing the challenges of insufficient and excessive editing has become a critical issue.
 To address the above limitations, we propose a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for AQA (As shown is Figure 1 (II)).
 Specifically, ARE first prompts LLMs to generate long-form answers for the given question.
-Subsequently, ARE utilizes an instruction-tuned LLM, trained on a well-constructed fact decomposition dataset, to decompose the long-form answers into molecular clauses and then into atomic facts.
-These atomic facts are then used to search the evidences from a search engine.
+Subsequently, ARE utilizes an instruction-tuned LLM [^1] The fine-tuned atomic fact decomposition LLM can be found in https://github.com/zhichaoyan11/FIDES, trained on a well-constructed fact decomposition dataset, to decompose the long-form answers into molecular clauses and then into atomic facts.
+These atomic facts are then used to search the evidence from a search engine.
 Additionally, ARE employs an evidence verifier to classify the relationships between the evidence and the atomic facts into three categories: 1) supportive, 2) editing required, and 3) irrelevant.
@@
   • This paper designs an evaluation metric $Attr_{p}$ to accurately assess the precision and completeness of retrieved evidences, mitigating the proportion of invalid retrieved evidences.
-Figure 2.
-An example of the ARE process.
+Figure 2: An example of the ARE process.
 In Molecular-to-Atomic stage, an instruction-tuned fact decomposition LLM decomposes the long-form answer $X$ into molecular clauses and atomic facts.
@@
 The Attribution Report consists of all evidences $\{E_{1},...,E_{m}\}$, where $MF_{i}^{{}^{\prime}}$ and $e_{ij}^{{}^{\prime}}$ are the edited contents.
-2. Related Work
-2.1. Attributed Question Answering
+II Related Work
+II-A Attributed Question Answering
 From the perspective of retrieval timing, two notable trends have recently emerged: (1) Retrieval-then-Read (RTR) and (2) Post-hoc Retrieval.
 Trivedi et al.
-(trivedi2022interleaving, ) introduce IRCoT, a method that interleaves retrieval with steps in a Chain of Thought (CoT).
+[15] introduces IRCoT, a method that interleaves retrieval with steps in a Chain of Thought (CoT).
 This approach not only guides the retrieval process using CoT but also utilizes the retrieved results to enhance the CoT.
 Muller et al.
-(muller-etal-2023-evaluating, ) investigate attribution in cross-lingual question answering (QA).
-ALCE (gao2023enabling, ) employs various methods to integrate retrieved documents into large language models (LLMs) for answer generation.
+[16] investigates attribution in cross-lingual question answering.
+ALCE [7] employs various methods to integrate retrieved documents into LLMs for answer generation.
 Li et al.
-(li2023llatrieval, ) introduce a progressive selection of evidence using LLMs with a classification-based prompting template.
+[17] introduces a progressive selection of evidence using LLMs with a classification-based prompting template.
 Bohnet et al.
-(bohnet2022attributed, ) conduct an extensive evaluation of LLM attributions, finding that while the Retrieval-then-Read (RTR) approach performs well, it necessitates the comprehensive use of a traditional training set, thereby highlighting the potential of post-hoc retrieval.
-RARR (gao2023rarr, ) is the first framework to implement question decomposition-based retrieval followed by revision.
-Building on RARR, PURR presents an end-to-end editor for text revision (chen2023purr, ).
+[8] conducts an extensive evaluation of LLM attributions, finding that while the Retrieval-then-Read (RTR) approach performs well, it necessitates the comprehensive use of a traditional training set, thereby highlighting the potential of post-hoc retrieval.
+RARR [9] is the first framework to implement question decomposition-based retrieval followed by revision.
+Building on RARR, PURR presents an end-to-end editor for text revision [10].
 Kang et al.
-(kang2023mitigating, ) propose a combination of RTR and post-hoc retrieval strategies.
+[18] proposes a combination of RTR and post-hoc retrieval strategies.
 The limitations of question decomposition were discussed in the Introduction.
 Our approach addresses these limitations through molecular-to-atomic fact decomposition and atomic-to-molecular editing, resulting in significant performance improvements.
-2.2. Fact Decomposition
-The technology of decomposition has been shown to effectively address complex questions, particularly in various reasoning tasks (lin2023decomposing, ; Ye_2023, ; zhang2023reasoning, ) and claim verification tasks (lin-etal-2023-decomposing, ; chen2023complex, ; kim2023kggpt, ; wang2023hallucination, ; kamoi-etal-2023-wice, ).
-While fact decomposition can accurately generate sub-facts that represent the original answer, it struggles to aggregate evidence for sentences based on the decomposition results.
+II-B Question or Fact Decomposition
+The technology of decomposition has been shown to effectively address complex questions, particularly in various Question Answering (QA) tasks (e.g., Multi Hop Question Answering (MHQA)) [19, 20, 21, 22] and claim verification tasks [23, 24, 25, 26, 27].
+Under the MHQA task, [28, 29, 30, 31, 32] typically decompose the original question or introduce retrieval augment generation methods to mitigate the ambiguity of questions and the model parameter knowledge limitations, respectively.
+While fact decomposition can generate sub-facts that represent the original answer, it struggles to decompose long-form answers that have a complex semantic structure and aggregate evidence for sentences based on the decomposition results.
 This limitation becomes evident when retrieving evidence by fact to support each sentence that contains more than one fact.
-2.3. Hallucination Detection and Revision
-Hallucination detection (chen2023hallucination, ; wang2023hallucination, ; kadavath2022language, ; mndler2023selfcontradictory, ; mishra2024fine, ) is a challenging yet essential task for improving the reliability of large language models (LLMs) in real-world scenarios.
+II-C Hallucination Detection and Revision
+Hallucination detection [33, 26, 34, 35, 36] is a challenging yet essential task for improving the reliability of LLMs in real-world scenarios.
 Similarly, Zheng et al.
-(ZLLP2024, ) introduced TrustScore, the first effective evaluation metric designed to assess the trustworthiness of LLM responses in a reference-free context.
-To tackle hallucinations, numerous model editing methods have been developed that do not involve updating the parameters of large language models (LLMs) (zheng-etal-2023-edit, ; gao2023rarr, ; HLTW+2023, ; dhuliawala2023chainofverification, ; wu2024updating, ; song2024knowledge, ).
+[37] introduced TrustScore, the first effective evaluation metric designed to assess the trustworthiness of LLMs responses in a reference-free context.
+Yan et al.
+[38] proposed CRAG, which trains an evidence evaluator model to assess the quality of retrieved evidence and classify retrieval actions as “Correct,” “Incorrect,” or “Ambiguous”.
+Different from our framework, CRAG aims to improve retrieval robustness by evaluating and refining retrieved evidence before generation and focuses on improving evidence quality rather than ensuring attribution and alignment with source material.
+To tackle hallucinations, numerous model editing methods have been developed that do not involve updating the parameters of LLMs [39, 9, 40, 12, 14, 41].
 In contrast to our approach, most of these methods focus exclusively on either hallucination detection or revision.
-3. Preliminary
+III Preliminary
 Attributed Question Answering (AQA) is a task which provides both a trustworthy answer and a reliable attribution report for a given question.
@@
 $X,\ A\longleftarrow\mathcal{M}_{\text{AQA}}(q,\ D),$
-where $X$ represents the long-form answer, $A$ is an attribution report consisting of of a collection of evidence, and $\mathcal{M}_{\text{AQA}}$ is a model used for the AQA task.
+where $X$ represents the long-form answer, $A$ is an attribution report consisting of a collection of evidence, and $\mathcal{M}_{\text{AQA}}$ is a model used for the AQA task.
 Post-hoc Retrieval-based AQA aims to enhance the reliability of LLM-generated long-form answers by further incorporating evidence retrieval, fact verification, and factual editing.
@@
 Symbol definitions.
-The key symbols used in our framework are depicted as follows: $MF_{i}$ represents a molecular clause within the long-form answer $X$; $AF_{ij}$ denotes the atomic fact of $MF_{i}$; $e_{ij}$ corresponds to the evidence for each atomic fact $AF_{ij}$; $E_{i}$ represents the aggregated evidence for the molecular clause $MF_{i}$, compiled from multiple evidence $e_{ij}$.
-Entropy measures the uncertainty or information content associated with random variables.
-The higher the entropy, the more uncertain or random the information is (ash2012information, ).
-It can be calculated as:
+The key symbols used in our framework are defined as follows: $MF_{i}$ denotes a molecular clause in the long-form answer $X$ that contains at least one fact subject to further decomposition; $AF_{ij}$ represents an atomic fact derived from $MF_{i}$ that cannot be further decomposed; $e_{ij}$ refers to the evidence retrieved for each atomic fact $AF_{ij}$; and $E_{i}$ denotes the aggregated evidence for the molecular clause $MF_{i}$, compiled from the corresponding evidence pieces $e_{ij}$.
+Entropy measures the uncertainty or information content associated with random variables. The higher the entropy, the more uncertain or random the information is [42]. It can be calculated as:
 $H(X)=-\sum_{i=1}^{n}P(x_{i})\log P(x_{i}),$
@@
 As demonstrated by Passalis et al.
-(7439840, ) that minimizing entropy can enhance performance in retrieval tasks.
+[43] that minimizing entropy can enhance performance in retrieval tasks.
 Long-form answers, characterized by greater complexity, have higher entropy, which makes it more challenging to retrieve relevant evidences.
-Traditional post-hoc retrieval based AQA methods use LLMs to generate sub-questions directly, which may contain multiple meanings and uncertainties.
+Traditional post-hoc retrieval based AQA methods use LLMs to generate sub-questions directly, which may contain multiple meanings, uncertainties and lose some key entities.
 Thus, these sub-questions often have high entropy, leading to retrieving ambiguous or irrelevant evidence (As shown in Figure 1 (I)).
-In contrast, fact decomposition preserves higher certainty and lowers entropy, thereby enhancing evidence retrieval (ash2012information, ).
-The experimental results in Section 5.5 verify our hypothesis by demonstrating that atomic-fact based retrieval outperforms sub-question based retrieval.
-Figure 3.
-The process of dataset construction and instruction tuning.
+In contrast, fact decomposition preserves higher certainty and lowers entropy, thereby enhancing evidence retrieval [42].
+To support this hypothesis, we begin with the following assumptions:
+1) $X$ is the long-form answer, $\mathrm{h}$ is a deterministic extractor that maps each $X$ to a set of entities: $E=\mathrm{h}(X)$, hence by the definition of conditional entropy for a deterministic function, $H(E|X)=0$.
+2) The support set for atomic fact $AF$ is $S_{AF}(X)=\{af:P(AF=af|X)>0\}$ and for sub-question $S_{Q}(X)={q:P(Q=q|X)>0}$.
+To fully cover the same triple fact $(e_{h},r,e_{t})$, the LLMs must generate at least two sub-questions (one for head entity and one for tail).
+Hence
+$\mid S_{Q}(X)\mid\geq 2\mid S_{AF}(X)\mid$ (1)
+3) From atomic fact set $AF$, $E$ can be recovered, the entropy $H(E|AF)=0$, similarly, from sub-question set $Q$, the entropy $H(E|Q)=0$.
+The theory of maximum entropy: For a length of support set $\mid\mathrm{supp(P)\mid}=M$, the upper bound of entropy is $H(P)\leq\mathrm{log}M$.
+The generation process of the support set is considered without any bias, the maximum entropy is uniform distribution.
+Therefore, we adopt $H(AF|X)\approx\mathrm{log}\mid S_{AF}(X)\mid$ and $H(Q|X)\approx\mathrm{log}\mid S_{Q}(X)\mid$.
+From eq 1,
+$\mathrm{log}\mid S_{Q}\mid\geq\mathrm{log}(2\mid S_{AF}\mid)=\mathrm{log}2+\mathrm{log}\mid S_{AF}\mid,$ (2)
+consider the upper bound of entropy,
+$H(Q|X)\geq\mathrm{log}2+\mathrm{log}\mid S_{AF}(X)\mid\geq\mathrm{log}2+H(AF|X)$ (3)
+$H(Q|X)-H(AF|X)\geq\mathrm{log}2>0$ (4)
+$\Rightarrow H(AF|X)<H(Q|X).$
+The theory of chain rule of entropy: Given the two random variables $(A|B)$, the chain is:
+$H(A,B)=H(A)+H(B|A)=H(B)+H(A|B).$ (5)
+$\mathrm{For}\ AF\mathrm{:}\ H(E,AF)$ $=H(AF)+H(E|AF)$
+$=H(AF)+0$
+$=H(E)+H(AF|E)$
+$\Rightarrow H(AF)=H(E)+H(AF|E).$
+For $Q$,
+$H(E,Q)$ $=H(Q)+H(E|Q)$
+$=H(Q)+0$
+$=H(E)+H(Q|E)$
+$\Rightarrow H(Q)=H(E)+H(Q|E).$
+Due to $H(E|X)=0$, $H(AF)=H(E)+H(AF|X)$, $H(Q)=H(E)+H(Q|X)$. The relation of $H(Q)$ and $H(AF)$ is
+$H(Q)-H(AF)$
+$=[H(E)+H(Q|X)]-[H(E)+H(AF|X)]$
+$=H(Q|X)-H(AF|X)\geq\mathrm{log}2\Rightarrow H(Q)>H(AF).$
+The entropy of atomic fact is lower than sub-question.
+The experimental results in Section V-H also verify our hypothesis by demonstrating that atomic-fact based retrieval outperforms sub-question based retrieval.
+Figure 3: The process of dataset construction and instruction tuning.
 We first extract the neighbors of the corresponding entity from Wikidata, and then utilize an LLM to convert each triple into a single atomic fact.
@@
 After two prompts, we obtain the dataset for instruction-tuning with LoRA.
-4. Approach
+IV Approach
 In this section, a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework is proposed for AQA task.
-Specifically, our ARE framework contains three stages: 1) Molecular-to-Atomic Fact Decomposition generally decomposes the generated long-form answer into molecular clauses and atomic facts by an instruction-tuned LLM; 2) Atomic Fact based Evidence Retrieval employs a search engine to retrieve evidence and selects the most similar evidence using a reranking model; 3) LLM-based Evidence Verifing and Editing guides LLMs through a verified result to either expand atomic facts for further re-retrieval or edit atomic facts with retrieved evidence.
-4.1. Molecular-to-Atomic Fact Decomposition
+Specifically, our ARE framework contains three stages: 1) Molecular-to-Atomic Fact Decomposition generally decomposes the generated long-form answer into molecular clauses and atomic facts by an instruction-tuned LLM; 2) Atomic Fact based Evidence Retrieval employs a search engine to retrieve evidence and selects the most similar evidence using a reranking model; 3) LLM-based Evidence Verifying and Editing guides LLMs through a verified result to either expand atomic facts for further re-retrieval or edit atomic facts with retrieved evidence.
+IV-A Molecular-to-Atomic Fact Decomposition
 This stage introduces the process that prompts LLMs to generate a long-form answer, and utilizes the instruction-tuned LLMs fine-tuned on a carefully constructed dataset to decompose the long-form answer into molecular clauses and atomic facts.
-4.1.1. Long-form answer generation
+IV-A1 Long-form answer generation
 LLMs are trained on the large-scale corpora which endow them with powerful generative capabilities.
-Long-form answers are generated by feeding few-shot prompting template and a specific question to LLMs.
+Long-form answers are generated by feeding a few-shot prompting template and a specific question to LLMs.
 Formally, the whole process can be represented as:
@@
 Tom Brady is an American football quarterback who has won six Super Bowl championships.
-4.1.2. Fact decomposition
-Fact decomposition aims to decompose long-form answers characterized by complex logical structures into molecular clauses and multiple atomic facts.
-Specifically, our ARE framework constructs a molecular-to-atomic fact decomposition dataset for instruction-tuning LLM, and this process enhances LLM’s ability to effectively manage the intricacies of long-form answers.
+IV-A2 Fact decomposition
+Fact decomposition aims to decompose long-form answers characterized by complex logical structures into molecular clauses and multiple atomic facts through the instruction-tuning LLM.
+Specifically, the fact decomposition module consists of two key components: data construction and the instruction-tuning of the Molecule-to-Atomic LLM.
 Dataset Construction.
-To fine-tune the molecule-to-atomic fact decomposition LLM, ARE constructs an instruct-tuning dataset that converts triples into texts containing molecule clauses and atomic facts derived from Knowledge Graphs (KGs) using LLMs.
+To fine-tune the molecule-to-atomic fact decomposition LLM, ARE employs LLMs to construct an instruct-tuning dataset that converts triples into texts containing molecule clauses and atomic facts derived from Knowledge Graphs (KGs) [44, 45].
 As depicted in Figure 3, the dataset construction process is divided into two steps: data collection and data processing with LLMs.
-Step1: ARE first randomly selects the entities from the Knowledge Graph Question Answering (KGQA) dataset (gu2021beyond, ) and then extracts the triplets of given entities from Wikidata [^1] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=.
+Step1: ARE first randomly selects the entities from the Knowledge Graph Question Answering (KGQA) dataset [46] and then extracts the triplets of given entities from Wikidata [^2] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=.
 As some properties or objects in the triplets may be useless (e.g., JPEG, code, and ISSN), they are thus further filtered by heuristic rules in the extracting process.
-Details of the rules can be found in Appendix C.4.
 Step2: ARE utilizes LLMs to randomly transform several triplets into atomic facts $\{AF_{i1},...,AF_{ik}\}$.
-The atomic facts are fed into LLMs to generate a response in the JSON format: {“Generated content”: $S$, “Molecular clauses”: $MF_{i}$, “Atomic facts”: $AF_{ij}$}, where $S$ is text consisting of $n$ molecular clauses, $MF_{i}$ is $i$-th molecular clause, $i\in\{1,...,m\}$, and $AF_{ij}$ is the $j$-th atomic fact of the $MF_{i}$, $j\in\{1,...,k\}$.
-Through the above process, an instruction-tuning dataset can be obtained, which contains 7138 samples for training and 1000 for evaluation. An example data can be found in Appendix A.2.
+The atomic facts are fed into LLMs to generate a response in the Json format: {“Generated content”: $S$, “Molecular clauses”: $MF_{i}$, “Atomic facts”: $AF_{ij}$}, where $S$ is text consisting of $m$ molecular clauses, $MF_{i}$ is $i$-th molecular clause, $i\in\{1,...,m\}$, and $AF_{ij}$ is the $j$-th atomic fact of the $MF_{i}$, $j\in\{1,...,k\}$.To retain the overall coherence and logical flow of the original long form answer, we format the molecular clauses and atomic facts into a Json structure: {“output”:[$MF_{i}$$AF_{ij}$...}…]}, which enhances the relationship among $MF_{i}$, $AF_{ij}$ and $X$.
+Through the above process, an instruction-tuning dataset can be obtained, which contains 7138 samples for training and 1000 for evaluation.
 Instruction-tuning the Molecule-to-Atomic LLM.
-For the fine-tuning stage, ARE employs the Low-Rank Adaptation (LoRA) technique (hu2022lora, ) for instruction-tuning the Llama3-8B-Instruct LLM to reduce computational complexity and time consumption.
+For the fine-tuning stage, ARE employs the Low-Rank Adaptation (LoRA) technique [47] for instruction-tuning the Llama3-8B-Instruct LLM to reduce computational complexity and time consumption.
 Specifically, LoRA works by freezing parameters $\theta_{0}$ of the pre-trained model, while adding trainable parameters $\Delta\theta_{0}$ that can be expressed as the product of two low-rank matrices:
-(1) $\Delta\theta_{0}=\mathbf{BA},$
+$\Delta\theta_{0}=\mathbf{BA},$ (6)
 where $\mathbf{B}\in\mathbb{R}^{d\times r}$, $\mathbf{A}\in\mathbb{R}^{r\times k}$, $r\ll\mathrm{min}(d,k)$.
 After obtaining the instruction-tuned Molecule-to-Atomic LLM, ARE decomposes the long-form answer “$X$: The player with the most Super Bowl rings is Tom Brady.
-Tom Brady is an American football quarterback who has won six Super Bowl championships.” into molecular clauses “$MF_{1}$: The player with the most Super Bowl rings is Tom Brady.” and “$MF_{2}$: Tom Brady is an American football quarterback who has won six Super Bowl championships.”, $MF_{1}$ only has one atomic fact $AF_{11}$, $MF_{2}$ can be further decomposed into two atomic facts “$AF_{21}$: Tom Brady is an American football quarterback.” and “$AF_{22}$: Tom Brady has won six Super Bowl championships.”.
-4.2. Atomic Fact based Evidence Retrieval
+Tom Brady is an American football quarterback who has won six Super Bowl championships.” into molecular clauses $MF_{1}$: The player with the most Super Bowl rings is Tom Brady.” and $MF_{2}$: Tom Brady is an American football quarterback who has won six Super Bowl championships.”, $MF_{1}$ only has one atomic fact $AF_{11}$, $MF_{2}$ can be further decomposed into two atomic facts $AF_{21}$: Tom Brady is an American football quarterback.” and $AF_{22}$: Tom Brady has won six Super Bowl championships.”.
+IV-B Atomic Fact based Evidence Retrieval
 Atomic fact based evidence retrieval aims to employ search engines [^3] We select Google Search as knowledge source, accessible via https://customsearch.googleapis.com/customsearch/v1 to search evidences that support atomic facts.
@@
 ARE leverages search engines to search evidence from each atomic fact $AF_{ij}$, while uses the pre-trained Sentence-Bert model [^4] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2 to embed both the searched evidences $e_{c}$ and the atomic fact $AF_{ij}$ into a common vector space:
-(2) $\mathbf{e}_{c}^{z},\text{{AF}}_{ij}=\textrm{Sentence-Bert}(e_{c}^{z},AF_{ij}),$
-where $e_{c}^{z},z\in\{1,...,o\}$ is the $z$-th searched evidence in $e_{c}$. ARE calculates the relevance score $R$($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) through the cosine similarity function:
-(3) $R(\mathbf{e}_{c}^{z},\text{{AF}}_{ij})=\frac{\mathbf{e}_{c}^{z}\cdot\text{{AF}}_{ij}}{\parallel\mathbf{e}_{c}^{z}\parallel\parallel\text{{AF}}_{ij}\parallel},$
-[…]
-4.3. LLM-based Evidence Verifing and Editing
-Due to the fact that hallucinated knowledge probably exists in the long-form answers generated by LLMs, ARE utilizes the evidence verifing LLM to assess whether atomic facts need to be edited or re-retrieved through comparing with the retrieved evidence.
-The evidence verifier takes as input the atomic fact and its corresponding evidence, and outputs different statuses, including:
-(4) $\textrm{EV}(AF_{ij},e_{ij})=\left\{\begin{matrix}1,&\textit{supportive}\\ 2,&\textit{editing required}\\ 3,&\textit{irrelevant},\end{matrix}\right.$
-where $\text{EV}(\cdot,\cdot)$ denotes the evidence verifier, which determines the relationships between the atomic fact and evidence.
+$\mathbf{e}_{c}^{z},\text{{AF}}_{ij}=\textrm{Sentence-Bert}(e_{c}^{z},AF_{ij}),$ (7)
+where $e_{c}^{z},z\in\{1,...,o\}$ is the $z$-th searched evidence in $e_{c}$. ARE calculates the relevance score RR($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) through the cosine similarity function:
+$R(\mathbf{e}_{c}^{z},\text{{AF}}_{ij})=\frac{\mathbf{e}_{c}^{z}\cdot\text{{AF}}_{ij}}{\parallel\mathbf{e}_{c}^{z}\parallel\parallel\text{{AF}}_{ij}\parallel},$ (8)
+where $\cdot$” represents the inner product and $\parallel\cdot\parallel$ is the norm of the corresponding vector [3].
+Finally, ARE ranks all the evidences $\{e_{c}^{1},...,e_{c}^{o}\}$ of $AF_{ij}$ by relevance score RR($\mathbf{e}_{c}^{z}$, $\text{{AF}}_{ij}$) and selects the top as the most relevant evidence $e_{ij}$ for $AF_{ij}$.
+The search engine’s response may not always yield valid content, often due to poorly constructed queries or limitations within the search engine itself.
+IV-C LLM-based Evidence Verifying and Editing
+Given that hallucinated content may exist in long-form answers generated by LLMs, ARE employs an evidence-verifying LLM to assess the factual consistency of atomic facts.
+This verifier operates using a few-shot, prompt-based approach to determine whether an atomic fact should be retained, edited, or re-retrieved, based on a comparison with the corresponding retrieved evidence.
+The verifier takes as input an atomic fact along with its associated evidence and outputs one of several status labels, including:
+$\textrm{EV}(AF_{ij},e_{ij})=\left\{\begin{matrix}1,&\textit{supportive}\\ 2,&\textit{editing required}\\ 3,&\textit{irrelevant},\end{matrix}\right$. (9)
+where $\text{EV}(\cdot,\cdot)$ denotes the evidence verifying LLM, which determines the relationships between the atomic fact and evidence.
 Specifically, $1$ represents supportive, requiring no further action; $2$ means editing required, necessitating the revision of atomic facts; $3$ indicates irrelevant, which requires a new evidence retrieval.
-The prompt template of the evidence verifier can be found in Appendix A.3.
+To address potential misjudgments arising from LLM uncertainty, we implement an iterative verification mechanism.
+This process repeatedly validates results flagged with the statuses editing required and irrelevant, ensuring greater accuracy in the edited or re-retrieved content.
+The prompt template used by the evidence verifier is provided in Table XI.
+Algorithm 1: Atomic fact decomposition-based Retrieval and
+Editing framework.
+1:
+2:
+3:
+4:
+5:
+6:
+7:
+8:
+9:
+10:
+11:
+12:
+13:
+14:
+15:
+16:
+17:
+18:
+19:
+20:
+Input: The question q, the corpus of text passages D;
+$A\leftarrow$ {};
+$X$ $\leftarrow$ LLMs(q);
+MFMF, AFAF $\leftarrow$ Fact Decomposition($X$);
+for $MF_{i}\in MF$ do
+    for $AF_{ij}\in AF$ do
+      $e_{ij}$ $\leftarrow$ Evidence Retrieval($AF_{ij}$, $D$);
+      for iteration=0: 4 do
+          status $\leftarrow$ Evidence Verifying($AF_{ij}$, $e_{ij}$);
+          if status == 1 do
+            $AF[j]\leftarrow AF_{ij}$
+            break
+          if status == 2 do
+            $AF_{ij}$ $\leftarrow$ Editing($AF_{ij}$, $e_{ij}$)
+          if status == 3 do
+             Expanded $AF_{ij}$ $\leftarrow$ Fact Expansion($AF_{ij}$)
+             $e_{ij}\leftarrow$ Evidence Retrieval(Expanded $AF_{ij}$, $D$);
+    AA\ $\cup$ Aggregate($e_{ij}$)
+$X^{\prime}\leftarrow$ Backtrack(AFAF, MFMF)
+output: The revised answer $X^{\prime}$ and the attribution report AA.
 Figure 2 illustrates the process.
@@
 $AF_{22}$ needs to be edited based on evidence $e_{22}$.
-For irrelevant, ARE utilizes the fact expanding LLM to expand the atomic fact $AF_{21}$.
-The expanded fact contains more complete information, enabling more effective retrieval of relevant evidence.
-For example, ”Tom Brady” is expanded to ”Thomas Edward Patrick Brady Jr.”.
-The newly retrieved evidence will be re-verified by the evidence verifier, and this process will be repeated.
-Details of the prompt template and the effectiveness of fact expansion from the perspective of the information theory can be found in Appendix A.3 and Appendix B.
-For editing required, ARE designs prompts for LLMs based on in-context learning and chain-of-thought prompting techniques  (zheng-etal-2023-edit, ; gao2023rarr, ), guiding the LLMs to revise the atomic fact $AF_{22}$ using the retrieved evidence $e_{22}$.
+For irrelevant, since LLMs can provide relevant information to guide retrieval systems [48], ARE utilizes the fact expansion LLM to expand the atomic fact $AF_{21}$.
+The expanded fact contains more complete contextual information, enabling a more effective retrieval of relevant evidence.
+For example, “Tom Brady” is expanded to “Thomas Edward Patrick Brady Jr.” which is a more complete entity.
+The newly retrieved evidence $e_{ij}^{\prime}$ will be re-verified by the verifier, and this process will be repeated.
+The details of the prompt template can be found in Table XI.
+For editing required, ARE designs prompts for LLMs based on in-context learning and chain-of-thought prompting techniques  [39, 9], guiding LLMs to revise the atomic fact $AF_{22}$ using the retrieved evidence $e_{22}$.
 Since the edits are made at the atomic fact level rather than revising the entire answer $X$, ARE can precisely adjust the necessary details, thus minimizing unnecessary modification.
-The detailed editing instructions are shown in Appendix A.4.
-The edited atomic facts are re-verified and backtracked to their original positions within the molecular clauses, forming the final revised long-form answer $X^{\prime}$ along with the other molecular clauses that do not require editing.
-To further obtain the attribution report $A$, all evidences $e_{ij}$ are aggregated into a sequence $E_{i}$ to support the $MF_{i}$.
+Detailed editing instructions are shown in Table XI.
+The edited atomic facts are re-verified and reintegrated into their original positions within the molecular clauses, producing the final revised long-form answer $X^{\prime}$.
+This process also preserves molecular clauses that did not require editing.
+During backtracking, both molecular and atomic facts are provided in Json format alongside the original long-form answer as a reference.
+This approach follows the instruction-tuning dataset construction methodology, enabling the LLM to effectively manage interdependencies among atomic facts.
+The detailed backtracking prompt is presented in Table XI.
+To further obtain the attribution report $A$, all evidences $e_{ij}$ are aggregated into a sequence $E_{i}$ to support $MF_{i}$.
 Since each $AF_{ij}$ is derived from the decomposition of $MF_{i}$, there may be overlaps among the evidence $e_{ij}$.
@@
 Ultimately, $A$ is compiled as $\{E_{1},...,E_{i},...,E_{m}\}$.
-[…]
-Methods GPT-3.5 Llama3-70B Llama2-70B
-[…]
+To intuitively demonstrate the entire procedure of our proposed method, we have summarized the process in Algorithm 1.
+TABLE I: Evaluation results on API-based GPT-3.5 and GPT-4o-mini.
+When evaluating different LLMs, all methods also use the corresponding LLM for construction.
+$Attr_{r}$ and $Attr_{p}$ are used to evaluate the attribution, while Pres is used to assess the Preservation after editing.
+$F1_{PP}$ and $F1_{RP}$ are the harmonic means of the $Attr_{p}$ and Pres, and the $Attr_{r}$ and Pres metrics, respectively.
+Additionally, Average $\mathrm{F1_{PP}}$ (Ave-$\mathrm{F1_{PP}}$) and $\mathrm{F1_{RP}}$ (Ave-$\mathrm{F1_{RP}}$) are used to provide a comprehensive overview for each row.
+Note that the RARR paper does not provide access to the test datasets; thus, the reported results are reproduced from the publication.
+Bold indicates the best performance, while underlining indicates the second-best performance.
+Methods GPT-3.5 GPT-4o-mini
+$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $\mathrm{F1_{PP}} Ave$-$\mathrm{F1_{RP}}$
 NQ
-DRQA 0.424 0.647 - - - 0.382 0.620 - - - 0.483 0.522 - - - - -
-EFEC 0.598 0.042 0.762 0.080 0.670 0.490 0.058 0.717 0.107 0.582 0.357 0.058 0.719 0.107 0.477 0.098 0.576
-CCVER 0.624 0.066 0.928 0.123 0.747 0.597 0.071 0.921 0.132 0.724 0.423 0.068 0.813 0.126 0.539 0.127 0.670
-RARR 0.649 0.058 0.868 0.109 0.743 0.646 0.060 0.850 0.112 0.734 0.516 0.066 0.594 0.119 0.552 0.113 0.676
-ARE 0.670 0.756 0.910 0.826 0.772 0.682 0.739 0.898 0.811 0.759 0.584 0.682 0.926 0.785 0.716 0.807 0.749
+DRQA 0.424 0.647 - - - 0.265 0.576 - - - - -
+EFEC 0.598 0.042 0.762 0.080 0.670 0.492 0.053 0.688 0.098 0.546 0.089 0.608
+CCVER 0.624 0.066 0.928 0.123 0.747 0.563 0.051 0.931 0.097 0.702 0.110 0.728
+RARR 0.649 0.058 0.868 0.109 0.743 0.493 0.080 0.904 0.147 0.638 0.128 0.691
+ARE 0.670 0.756 0.910 0.826 0.772 0.623 0.722 0.902 0.802 0.737 0.814 0.755
 Mintaka
-DRQA 0.431 0.673 - - - 0.380 0.640 - - - 0.368 0.600 - - - - -
-EFEC 0.557 0.040 0.729 0.076 0.632 0.538 0.057 0.728 0.106 0.619 0.498 0.029 0.739 0.056 0.595 0.079 0.615
-CCVER 0.630 0.069 0.937 0.129 0.753 0.582 0.073 0.901 0.135 0.707 0.397 0.036 0.850 0.069 0.541 0.111 0.667
-RARR 0.646 0.060 0.850 0.112 0.734 0.651 0.065 0.829 0.121 0.729 0.543 0.058 0.679 0.107 0.603 0.113 0.689
-ARE 0.716 0.807 0.914 0.857 0.803 0.712 0.767 0.887 0.823 0.790 0.631 0.706 0.940 0.806 0.755 0.829 0.783
+DRQA 0.431 0.673 - - - 0.251 0.544 - - - - -
+EFEC 0.557 0.040 0.729 0.076 0.632 0.497 0.049 0.698 0.092 0.581 0.084 0.665
+CCVER 0.630 0.069 0.937 0.129 0.753 0.634 0.045 0.933 0.086 0.755 0.108 0.754
+RARR 0.646 0.060 0.850 0.112 0.734 0.508 0.058 0.899 0.109 0.649 0.111 0.692
+ARE 0.716 0.807 0.914 0.857 0.803 0.634 0.755 0.881 0.813 0.737 0.835 0.770
 StrategyQA
-DRQA 0.237 0.490 - - - 0.237 0.379 - - - 0.234 0.467 - - - - -
-EFEC 0.354 0.049 0.716 0.092 0.474 0.319 0.051 0.666 0.095 0.432 0.361 0.031 0.721 0.059 0.481 0.082 0.462
-CCVER 0.372 0.047 0.932 0.089 0.532 0.435 0.063 0.917 0.118 0.590 0.323 0.058 0.854 0.109 0.468 0.105 0.530
-RARR 0.356 0.097 0.862 0.174 0.504 0.449 0.073 0.846 0.134 0.586 0.412 0.057 0.604 0.104 0.490 0.138 0.527
-ARE 0.463 0.559 0.899 0.689 0.611 0.474 0.502 0.907 0.646 0.623 0.484 0.533 0.912 0.673 0.633 0.667 0.622
+DRQA 0.237 0.490 - - - 0.080 0.222 - - - - -
+EFEC 0.354 0.049 0.716 0.092 0.474 0.330 0.071 0.638 0.128 0.435 0.110 0.455
+CCVER 0.372 0.047 0.932 0.089 0.532 0.341 0.056 0.961 0.106 0.503 0.098 0.518
+RARR 0.356 0.097 0.862 0.174 0.504 0.292 0.079 0.916 0.145 0.443 0.160 0.474
+ARE 0.463 0.559 0.899 0.689 0.611 0.413 0.469 0.904 0.618 0.567 0.654 0.589
 ExpertQA
-DRQA 0.127 0.283 - - - 0.131 0.326 - - - 0.147 0.319 - - - - -
-EFEC 0.343 0.071 0.686 0.129 0.457 0.356 0.076 0.698 0.137 0.472 0.357 0.058 0.719 0.107 0.477 0.124 0.469
-CCVER 0.292 0.071 0.967 0.132 0.449 0.282 0.081 0.942 0.149 0.434 0.212 0.064 0.845 0.119 0.339 0.133 0.407
-RARR 0.340 0.078 0.904 0.144 0.494 0.400 0.084 0.851 0.153 0.544 0.353 0.074 0.606 0.132 0.446 0.143 0.495
-ARE 0.412 0.438 0.917 0.593 0.569 0.425 0.417 0.924 0.575 0.582 0.390 0.386 0.942 0.548 0.552 0.572 0.568
-5. Experiments
-In this section, we outline the experimental setups, present the experimental results, and provide a thorough experimental analysis.
-5.1. Evaluation Setups
-This section mainly introduces the datasets used and baseline methods for comparison, and the evaluation metrics employed.
+DRQA 0.127 0.283 - - - 0.122 0.308 - - - - -
+EFEC 0.343 0.071 0.686 0.129 0.457 0.297 0.075 0.635 0.134 0.405 0.132 0.431
+CCVER 0.292 0.071 0.967 0.132 0.449 0.310 0.062 0.980 0.117 0.471 0.221 0.460
+RARR 0.340 0.078 0.904 0.144 0.494 0.269 0.064 0.944 0.120 0.419 0.132 0.457
+ARE 0.412 0.438 0.917 0.593 0.569 0.387 0.414 0.914 0.570 0.544 0.582 0.557
+TABLE II: Evaluation results on the open-source Llama3-70B and Llama2-70B.
+The metrics and other experimental settings are the same as those used in API-based evaluation.
+Bold indicates the best performance, while underlining indicates the second-best performance.
+Methods Llama3-70B Llama2-70B
+$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $\mathrm{F1_{PP}} Ave$-$\mathrm{F1_{RP}}$
+NQ
+DRQA 0.382 0.620 - - - 0.483 0.522 - - - - -
+EFEC 0.490 0.058 0.717 0.107 0.582 0.357 0.058 0.719 0.107 0.477 0.107 0.530
+CCVER 0.597 0.071 0.921 0.132 0.724 0.423 0.068 0.813 0.126 0.539 0.129 0.632
+RARR 0.646 0.060 0.850 0.112 0.734 0.516 0.066 0.594 0.119 0.552 0.115 0.643
+ARE 0.682 0.739 0.898 0.811 0.759 0.584 0.682 0.926 0.785 0.716 0.798 0.738
+Mintaka
+DRQA 0.380 0.640 - - - 0.368 0.600 - - - - -
+EFEC 0.538 0.057 0.728 0.106 0.619 0.498 0.029 0.739 0.056 0.595 0.081 0.607
+CCVER 0.582 0.073 0.901 0.135 0.707 0.397 0.036 0.850 0.069 0.541 0.102 0.624
+RARR 0.651 0.065 0.829 0.121 0.729 0.543 0.058 0.679 0.107 0.603 0.114 0.666
+ARE 0.712 0.767 0.887 0.823 0.790 0.631 0.706 0.940 0.806 0.755 0.815 0.773
+StrategyQA
+DRQA 0.237 0.379 - - - 0.234 0.467 - - - - -
+EFEC 0.319 0.051 0.666 0.095 0.432 0.361 0.031 0.721 0.059 0.481 0.077 0.457
+CCVER 0.435 0.063 0.917 0.118 0.590 0.323 0.058 0.854 0.109 0.468 0.113 0.529
+RARR 0.449 0.073 0.846 0.134 0.586 0.412 0.057 0.604 0.104 0.490 0.119 0.538
+ARE 0.474 0.502 0.907 0.646 0.623 0.484 0.533 0.912 0.673 0.633 0.660 0.628
+ExpertQA
+DRQA 0.131 0.326 - - - 0.147 0.319 - - - - -
+EFEC 0.356 0.076 0.698 0.137 0.472 0.357 0.058 0.719 0.107 0.477 0.122 0.475
+CCVER 0.282 0.081 0.942 0.149 0.434 0.212 0.064 0.845 0.119 0.339 0.134 0.387
+RARR 0.400 0.084 0.851 0.153 0.544 0.353 0.074 0.606 0.132 0.446 0.142 0.495
+ARE 0.425 0.417 0.924 0.575 0.582 0.390 0.386 0.942 0.548 0.552 0.561 0.567
+V Experiments
+V-A Evaluation Setups
 Dataset.
-We perform extensive experiments on three Question Answering (QA) datasets: Natural Questions (NQ) (kwiatkowskinatural, ), Mintaka (sen2022mintaka, ), and StrategyQA (geva2021strategyqa, ), as well as the AQA dataset ExpertQA.
+We perform extensive experiments on three widely-used Question Answering (QA) datasets, as well as the AQA dataset, i.e., ExpertQA.
+Natural Questions (NQ) [49] is a dataset for answering questions developed by Google Search and is widely used to evaluate machine reading comprehension, information retrieval, and other tasks.
+StrategyQA [50] requires models to provide the reasoning process to come up with correct answers.
+Because, the correct answers are typically based on different paragraphs from multiple documents.
+Mintaka [51] is a complex multilingual dataset featuring question types such as superlative and multi-hop, which is aknowledge-intensive dataset and widely used to evaluate end-to-end question answering models.
+ExpertQA [52] encompasses 7 question types across 32 domains and is specifically designed to evaluate attribution-based question answering models.
 These datasets are used to evaluate the attribution ability in knowledge-intensive QA task.
-Following the standard dataset settings (gao2023rarr, ), we randomly select 150 samples from NQ, Mintaka and StrategyQA as test datasets.
+Following the standard dataset settings [9], we randomly select 150 samples from NQ, Mintaka, and StrategyQA as test datasets.
 For ExpertQA, we use the entire provided test set.
-Details of the datasets can be found in Appendix C.1.
 Baselines.
-The proposed ARE is compared with four post-hoc retrieval based baselines, including: EFEC [^4] https://github.com/j6mes/acl2021-factual-error-correction (thorne-vlachos-2021-evidence, ), DRQA (bohnet2022attributed, ), RARR [^5] https://github.com/anthonywchen/RARR (gao2023rarr, ) and CCVER [^6] https://github.com/jifan-chen/Fact-checking-via-Raw-Evidence (chen-etal-2024-complex, ).
-More details about the baselines can be found in Appendix C.2.
-Metrics. We assess the attribution and editing through several metrics: $Attr_{r}$ is a molecular-level attribution metric, which measures the recall of retrieved evidences (gao2023rarr, ):
-(5) $Attr_{r}(X,A)=\underset{MF_{i}\in X}{\textrm{avg}}\ \underset{E_{i}\in A}{\textrm{max}}\ \mathrm{NLI}(E_{i},MF_{i}),$
+The proposed ARE is compared with four post-hoc retrieval based baselines, including: DRQA [8] is a straightforward attribution model that concatenates the question and answer into a single query before retrieval.EFEC [^5] https://github.com/j6mes/acl2021-factual-error-correction [53] is an editor built on the T5 model and modifies a claim by incorporating evidence.
+RARR [^6] https://github.com/anthonywchen/RARR [9] first generates a series of sub-questions based on the long-form answer and retrieve evidence using these sub-questions as queries.
+The evidence will be used to verify the claim and further revise it by LLMs.
+CCVER [^7] https://github.com/jifan-chen/Fact-checking-via-Raw-Evidence [54] is a model that focusses on fact verification, it uses a few-shot approach to prompt LLMs to generate “Yes or No” type sub-questions for a given claim, which are then used as queries to retrieve evidence.
+To ensure a fair comparison across different baselines, we use the same retrieval settings for EFEC as we did for RARR, and we employ RARR’s editing module for CCVER, retaining the sub-questions generation prompt of CCVER to examine the impact of generating various types of questions on attribution performance.
+LLMs.
+We evaluated four LLMs: two open-source models (Llama2-70B-Chat, Llama3-70B-Instruct) and two API-based models (GPT-3.5-Turbo-1106, GPT-4o-mini-2024-07-18).
+For each specific task (e.g., long-form answer generation), the same model version was used across experiments, with variations only in prompt templates tailored to the task.
+The only exception is our instruction-tuned LLama3-8B-Instruct model, which is specifically fine-tuned for the fact decomposition task.
+Metrics. We assess the attribution and editing through several metrics: $Attr_{r}$ is a molecular-level attribution metric, which measures the recall of retrieved evidences [9]:
+$Attr_{r}(X,A)=\underset{MF_{i}\in X}{\textrm{avg}}\ \underset{E_{i}\in A}{\textrm{max}}\ \mathrm{NLI}(E_{i},MF_{i}),$ (10)
 where, $MF_{i}\in X$ is a molecular clause and $E_{i}\in A$ is the evidence for $MF_{i}$, “max” selects the highest entailment score among all evidences, and “avg” calculates the average score of all evidence, $\mathrm{NLI}(E_{i},MF_{i})$ [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture represents the model probability of $E_{i}$ entailing $MF_{i}$.
+The NLI model is based on T5-11B, which is trained on six datasets: MNLI, SNLI, FEVER, PAWS, SciTail, and VitaminC  [55, 56, 57, 58, 59, 60].
+The input format is: “premise: {evidence} hypothesis: {sentences}”.
+As suggested by [9], we use the probability whose predicting result is “1” as the entailment score in the metric $Attr_{r}$.
 However, $Attr_{r}$ focuses solely on evidence recall, neglecting the precision of the recalled evidence.
 This can result in high scores when the quantity of evidence is sufficiently large.
-In light of this, we propose $Attr_{p}$ to simultaneously assess the precision and completeness of the evidences by calculating the proportion of invalid evidence among all evidences, which is defined as:
-(6) $Attr_{p}=\frac{\sum_{i=1}^{m}\mathbb{I}({\textrm{NLI}_{bi}(E_{i},MF_{i})})}{m},$
-where $\mathbb{I}(\mathrm{condition})$ is 1 if the condition is true, 0 for other wise; $E_{i}\in A$ and $MF_{i}\in X$.
-If $E_{i}$ can not entail $MF_{i}$, it was defined as invalid evidence.
-As for $\mathrm{NLI}_{bi}(\cdot,\cdot)$ is a binary classification result based on TRUE model.
-Different from the $\mathrm{NLI}(\cdot,\cdot)$, $\mathrm{NLI}_{bi}(\cdot,\cdot)$ will return true only when the evidence supports the whole sentence, which makes $Attr_{p}$ more stricter than $Attr_{r}$.
-The proof of $Attr_{p}$ can be used to evaluate the precision is available in Appendix E.
-Preservation (gao2023rarr, ) generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$:
-(7) $Pres_{(X,X^{\prime})}=\textrm{max}(1-\frac{\textrm{Lev}(X,X^{\prime})}{\textrm{length}(X)},0),$
-where $Pres_{(X,X^{\prime})}$ equals 1 when $X$ is identical to $X^{\prime}$, indicating no changes. A value of 0 means $X$ and $X^{\prime}$ share no common words, reflecting complete divergence.
-To better compare with different baselines, $F1_{RP}$ (gao2023rarr, ; chen2023purr, ) and $F1_{PP}$ is proposed. $F1_{RP}$ and $F1_{PP}$ are calculated by the following equations:
-[…]
-More details of the metrics can be found in Appendix C.3.
-Table 2.
-The ablation experiment results on NQ, StrategyQA, and ExpertQA using GPT-3.5.
-“w/o” represents removal for the mentioned module.
-The ablation study on Mintaka is shown in Appendix C.2.
-We mark the better results in bolded.
-Methods NQ StragegyQA ExpertQA
-[…]
-w/o edit 0.648 0.715 - - - 0.442 0.53 - - - 0.394 0.417 - - -
-w/o atomic 0.451 0.619 0.9 0.734 0.608 0.225 0.331 0.918 0.487 0.362 0.168 0.308 0.943 0.464 0.285
-w/o molecular 0.713 0.785 0.585 0.670 0.643 0.514 0.609 0.598 0.603 0.553 0.425 0.449 0.561 0.499 0.484
-w/o re-retrieval 0.609 0.677 0.911 0.777 0.730 0.437 0.533 0.927 0.677 0.594 0.388 0.421 0.937 0.581 0.548
-ARE 0.69 0.737 0.91 0.814 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
-5.2. Experimental Results and Analysis
-The experimental results are displayed in Table 1, and the experimental analyses are listed as follows:
-(1) ARE demonstrates significant improvements in both molecular-level attribution evaluation metrics and intent preservation across four datasets for all LLMs.
-ARE achieves improvements of 68%, 71.6%, 52.9%, and 42.9% over the $Ave-\mathrm{F1_{PP}}$ metric across all datasets.
-Additionally, ARE also achieves the best performance in the $Ave-\mathrm{F1_{RP}}$ metric, with improvements of 7.3%, 9.4%, 9.2%, and 7.3% on GPT-3.5, Llama3-70B, and Llama2-70B [^8] We use GPT-3.5-turbo-1106 for reproducibility.
-The Llama series models can be downloaded from https://huggingface.co/meta-llama/Llama-2-70b-chat and https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct., respectively, demonstrating the superior effectiveness and generalization capabilities of our proposed ARE.
-(2) For attribution ability, ARE outperforms other baselines on $Attr_{r}$ and $Attr_{p}$.
-Especially in $Attr_{p}$, all existing methods are far below ARE, this is because $Attr_{p}$ is a strict metric that evaluates the completeness of evidence.
-As for Pres, although the CCVER gets a higher $Pres$ in GPT-3.5 and Llama3-70B, it performed poorly on Llama-2-70B.
-In contrast, ARE achieved the most stable intent preservation across all LLMs, even with the less effective Llama2-70B model, which struggles with intent preservation, particularly when using the RARR method.
-These results demonstrate that the atomic fact-base ARE can maximize the attribution score while preserving the original intent.
-(3) The sub-question generation methods perform similarly on NQ, Mintaka, and StrategyQA, but CCVER significantly underperforms compared to RARR on ExpertQA.
-This highlights the limitations of sub-question generation, which require carefully designed prompts.
+To this end, we propose $Attr_{p}$ to simultaneously assess the precision and completeness of the evidences by calculating the proportion of invalid evidence among all evidences:
+$Attr_{p}=\frac{\sum_{i=1}^{m}\mathbb{I}({\textrm{NLI}_{bi}(E_{i},MF_{i})})}{m},$ (11)
+where $\mathbb{I}(\mathrm{condition})$ is 1 if the condition is true, 0 otherwise; $E_{i}\in A$ and $MF_{i}\in X$.
+The aggregated evidence $E_{i}$ for a molecular clause $MF_{i}$ consists of multiple evidence snippets $e_{ij}$, each corresponding to an atomic fact $AF_{ij}$.
+We define $E_{i}$ as “invalid” if any individual snippet $e_{ij}$ fails to support its associated atomic fact.
+To assess this, we use $\mathrm{NLI}_{bi}(\cdot,\cdot)$, a binary classification model based on the TRUE framework.
+Unlike the probabilistic version $\mathrm{NLI}(\cdot,\cdot)$, $\mathrm{NLI}_{bi}(\cdot,\cdot)$ returns “1” only if the evidence fully supports the entire sentence.
+This makes $Attr_{p}$ a stricter metric than $Attr_{r}$, enabling it to better evaluate the completeness of the retrieved evidence.
+Preservation [9] generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$:
+$Pres_{(X,X^{\prime})}=\textrm{max}(1-\frac{\textrm{Lev}(X,X^{\prime})}{\textrm{length}(X)},0),$ (12)
+where $Pres_{(X,X^{\prime})}$ equals 1 when $X$ is identical to $X^{\prime}$, indicating no changes.
+A value of 0 means $X$ and $X^{\prime}$ do not share common words, reflecting a complete divergence.
+To better compare with different baselines, $F1_{RP}$ [9, 10] and $F1_{PP}$ are proposed. $F1_{RP}$ and $F1_{PP}$ are calculated by the following equations:
+$\mathrm{F1_{RP}=\frac{2*Attr_{r}*Pres}{Attr_{r}+Pres}},\quad\mathrm{F1_{PP}=\frac{2*Attr_{p}*Pres}{Attr_{p}+Pres}}.$ (13)
+TABLE III: The ablation experiment results on NQ, StrategyQA, Mintaka and ExpertQA using GPT-3.5. “w/o” represents removal for the mentioned module. We mark the better results in bolded.
+Method NQ StragegyQA
+$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
+w/o edit 0.648 0.715 - - - 0.442 0.53 - - -
+w/o atomic level 0.451 0.619 0.9 0.734 0.608 0.225 0.331 0.918 0.487 0.362
+w/o molecular level 0.713 0.785 0.585 0.670 0.643 0.514 0.609 0.598 0.603 0.553
+w/o re-retrieval 0.609 0.677 0.911 0.777 0.730 0.437 0.533 0.927 0.677 0.594
+ARE 0.69 0.737 0.91 0.814 0.772 0.463 0.559 0.899 0.689 0.611
+Mintaka ExpertQA
+w/o edit 0.672 0.768 - - - 0.404 0.417 - - -
+w/o atomic level 0.477 0.651 0.928 0.765 0.630 0.168 0.308 0.943 0.464 0.285
+w/o molecular level 0.725 0.832 0.548 0.661 0.624 0.425 0.449 0.561 0.499 0.484
+w/o re-retrieval 0.68 0.767 0.936 0.843 0.788 0.388 0.421 0.937 0.581 0.548
+ARE 0.716 0.807 0.914 0.857 0.803 0.412 0.438 0.917 0.593 0.569
+Figure 4: The visualization experiments on NQ and ExpertQA datasets.
+The dashed line represents the highest attribution scores of all methods before editing.
+The points represent the performance of various methods after editing.
+The contours display level curves for $\mathrm{F1_{RP}}$ (left) and $\mathrm{F1_{PP}}$ (right), where points on the same contour share the same value.
+The closer a point is to the upper right corner, the better its performance is represented.
+V-B Experimental Results and Analysis
+The experimental results are displayed in Table I and Table II, and the experimental analyses are listed as follows:
+(1) ARE demonstrates substantial improvements in both molecular-level attribution evaluation metrics and intent preservation across four datasets using various LLMs.
+For instance, on Mintaka with API-based LLMs, ARE improves Ave-$\mathrm{F1_{PP}}$ by 72.4% and Ave-$\mathrm{F1_{RP}}$ by 1.6% (Table I).
+With open-source LLMs (Table II), ARE also achieves state-of-the-art performance—for example, in NQ, Ave-$\mathrm{F1_{PP}}$ and Ave-$\mathrm{F1_{RP}}$ improve by 9.5% and 68.3%, respectively—demonstrating the strong effectiveness and generalizability of our approach.
+The relatively smaller improvements observed on ExpertQA can be attributed to: (1) its specialized questions that often require expert-level domain knowledge beyond standard open knowledge bases, and (2) its significantly longer answers, averaging over 20 tokens more than other datasets.
+(2) For attribution ability, ARE outperforms other baselines in $Attr_{r}$ and $Attr_{p}$.
+Especially in $Attr_{p}$, all existing methods are far below ARE.
+This is because CCVER and RARR retrieve a larger quantity of evidence, which is typically fragmented.
+Although fragmented evidence can be beneficial for $Attr_{r}$ due to its focus on entailment probability, $Attr_{p}$ is a stricter metric that evaluates the completeness of the evidence.
+These evidence cannot support the entire sentence, resulting in poor performance in $Attr_{p}$.
+As for Pres, although the CCVER gets a higher Pres in GPT-3.5 and Llama3-70B, it performed poorly on Llama2-70B.
+In contrast, ARE achieved the most stable Pres across all LLMs, even with the less effective Llama2-70B, particularly when using the RARR method.
+These results demonstrate that ARE can maximize the attribution score while preserving the original intent.
+(3) The sub-question generation methods perform similarly on NQ, Mintaka and StrategyQA, but CCVER significantly underperforms compared to RARR on ExpertQA.
+This highlights the limitations of sub-question generation, which requires carefully designed prompts.
 Conversely, ARE shows greater robustness and adaptability across different datasets.
-Further experimental results are available in Appendix D.
-Figure 4. The performance of the fact decomposition LLM at the molecular level and the atomic level for different number of iterations.
-5.3. Ablation Study
-In order to investigate the impact of key modules on experimental performance, we conduct a series of ablation experiments, and the corresponding results are presented in Table 2.
-Specifically, “w/o edit” means removing the editing module; “w/o atomic” represent removing atomic-level facts; “w/o molecular” represents removing molecular-level clause; “w/o re-retrieval” means removing the status of “irrelevant” in evidence verifier.
+V-C Ablation Study
+In order to investigate the impact of key modules on experimental performance, we conduct a series of ablation experiments, and the corresponding results are presented in Table III.
+Specifically, “w/o edit” means removing the editing module; “w/o atomic” represents removing atomic-level facts; “w/o molecular” represents removing molecular-level clause; “w/o re-retrieval” means removing the status of “irrelevant” in evidence verifier.
 (1) In the “w/o edit” setting, attribution scores decline significantly.
-For example, $Attr_{r}$ decreased by 4.2%, while $Attr_{p}$ dropped by 2.2% in NQ.
-These results underscore the importance of atomic fact editing, as modifying hallucinated content is crucial for further improving attribution accuracy.
-(2) The results of the “w/o atomic” setting show the atomic facts play an important role in attribution scores.
-Specifically, ARE obtains 23.9% in $Attr_{r}$ and 11.8% in $Attr_{p}$ improvements on the NQ dataset.
-This phenomenon indicate that atomic facts have higher certainty and lower entropy, making it possible to retrieve relevant evidence more effectively.
-(3) Although the “w/o molecular” setting performs better on $Attr_{r}$ and $Attr_{p}$, it shows poorer performance on the $pres$ metric.
+For example, $Attr_{r}$ decreased by 4.2%, while $Attr_{p}$ dropped by 2.2% on NQ.
+These results underscore the importance of atomic fact editing, as modifying hallucinated content is crucial for further improving the accuracy of attribution.
+(2) The results of the “w/o atomic” setting show that atomic facts play an important role in the attribution scores.
+Specifically, ARE obtains improvements of 23.
+9% in $Attr_{r}$ and 11.
+8% in $Attr_{p}$ on NQ dataset.
+This phenomenon indicates that atomic facts have higher certainty and lower entropy, making it possible to retrieve relevant evidence more effectively.
+(3) Although the “w/o molecular” setting performs better on $Attr_{r}$ and $Attr_{p}$, it shows poorer performance on the Pres metric (decreased by 32.5% on NQ, 30.1% on StrategyQA, 35.6% on ExpertQA).
 This may be attributed to the removal of molecular facts, which causes atomic facts to lose their correspondence with them, preventing accurate backtracking to their original positions for editing.
@@
 (4) The performance of “w/o re-retrieval” shows that the necessity of the “irrelevant” status in evidence verifier module.
-For example, the $Attr_{r}$ decreased by 8.1% and 6% in $Attr_{p}$ in the NQ dataset.
+For example, $Attr_{r}$ decreased by 8.1% and 6% in $Attr_{p}$ in the NQ dataset.
 This phenomenon demonstrates that expanding facts when they have irrelevant evidence and then re-retrieving evidence can effectively enhance the attribution capability.
-Table 3. Comparison of the proposed atomic fact-based retrieval with previous sub-question-based retrieval performance.
+TABLE IV: Comparison with two prompt-based variants of ARE on Mintaka dataset using GPT-3.5.
+Methods $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
+CCVER 0.630 0.069 0.937 0.129 0.753
+RARR 0.646 0.06 0.850 0.112 0.734
+Prompt-based two-stage ARE 0.688 0.77 0.858 0.730 0.764
+Prompt-based ARE 0.682 0.768 0.774 0.723 0.725
+ARE 0.716 0.807 0.914 0.857 0.803
+V-D Visualization Results of Attribution Scores and Preservation
+To provide a more intuitive comparison between existing methods and our proposed ARE, regarding the simultaneous achievement of maximum attribution scores and preservation of original intent, we conducted the visualization experiments illustrated in Figure 4.
+ARE is located on the right side of the dashed line, indicating that ARE has achieved effective editing.
+The contour lines indicate that the proposed ARE achieves the best performance in both the F1RP (F1RP both considers Attrr and Pres) and F1PP (F1RP simultaneously considers Attrp and Pres) metrics after editing.
+These phenomena demonstrate the robustness of the proposed ARE in maximizing the attribution scores and preserving the original intent.
+V-E Effectiveness of Instruction-tuning
+To demonstrate the effectiveness of our proposed instruction-tuned fact decomposition LLM, we further present two variants of ARE: 1) Prompt-based two-stage ARE: This method first prompts LLMs to generate molecular clauses and then uses them to generate atomic facts using prompts.
+2) Prompt-based ARE: This method only replaces the instruction-tuned fact decomposition LLM with the prompt.
+As illustrated in Table IV, ARE significantly outperforms two variants of ARE, the performance decreased by about 12% and 4% in F1PP and F1RP, respectively.
+Prompt-based ARE fails to capture the complex structure of long-form answers, and prompt-based two-stage ARE still performs less than ARE, although it simplifies the process of generating atomic facts by first generating molecular clauses.
+Furthermore, ARE gives better performance because it utilizes a fact decomposition LLM tuned on a well-constructed dataset that enables ARE to better capture the complex structure of long-form answers.
+V-F Experiments of Aggregating Evidence from Sub-questions
+The sub-question generation method (such as RARR) lacks a clear correspondence between questions and clauses, thus failing to provide complete evidence for molecular clauses.
+To demonstrate this, we employ our Molecular-to-Atomic fact decomposition LLM to first decompose the long-form answer into molecular clauses and then generate sub-questions for every molecular clause using GPT-3.5.
+Finally, the evidences retrieved through the sub-questions are aggregated according to the molecular clause.
+As shown in Table V, the performance in $Attr_{p}$ is obviously improved (from 0.109 to 0.766 on NQ) by aggregating the evidence according to the molecular clauses (the method of “RARR w/ aggregate”), showing that $Attr_{p}$ can evaluate the completeness.
+However, $Attr_{r}$ decreases, indicating that RARR retrieves irrelevant evidence that makes the aggregated evidence noisier.
+TABLE V: The experimental results of aggregating evidence retrieved from sub-questions into the corresponding molecular clauses.
+Method NQ StrategyQA ExpertQA
+$Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$ $Attr_{r}$ $Attr_{p}$ Pres $\mathrm{F1_{PP}}$ $\mathrm{F1_{RP}}$
+RARR 0.649 0.058 0.868 0.109 0.743 0.356 0.097 0.862 0.174 0.504 0.34 0.078 0.904 0.144 0.494
+RARR $w/$ aggregate 0.499 0.646 0.942 0.766 0.653 0.336 0.53 0.948 0.680 0.496 0.276 0.407 0.959 0.571 0.428
+ARE 0.67 0.756 0.91 0.826 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
+V-G The Performance of Fact Decomposition and Quality Analysis of Constructed Dataset
+To find the best performance of molecular-to-atomic fact decomposition LLM and explore how iterations affect fact decomposition performance, we perform a comparative experiment, and the results are illustrated in Figure 5.
+Specifically, we assess its performance at both the molecular and atomic levels, focusing on two key aspects: consistency and correctness.
+Consistency $n_{c}$ measures how closely the number of decomposed sentences aligns with that of gold sentences.
+Correctness $d_{correct}$ evaluates whether the decomposed sentences preserve the original sentence meaning.
+Figure 5: The performance of the fact decomposition LLM at the molecular level and the atomic level for different number of iterations.
+We test on a well-constructed evaluation dataset mentioned in IV-A to select the decomposition model used in ARE.
+As shown in Figure 5, the performance of $n_{c}$ at the molecular-level initially increases and then decreases as the number of epochs increases.
+At the atomic-level, $n_{c}$ shows an overall upward trend, stabilizing at epoch 4.
+For $d_{correct}$, performance at molecular-level also increases and then decreases as the number of epochs increases, peaking at epoch 3 with 0.972.
+However, at the atomic-level, it first decreases and then increases.
+In terms of consistency, the performance at the 3rd and 4th epochs is similar.
+However, for correctness, the score of the atomic-level at epoch 3 is the highest, and meanwhile the score of the molecular-level at epoch 3 and epoch 4 has little difference.
+Therefore, we select the LLM trained with 3 epochs.
+Although the fact decomposition LLM has achieved satisfactory performance, misaligned decomposition scenarios still exist.
+Since ARE is a multi-stage pipeline architecture, such misaligned decomposition may lead to error propagation.
+This issue represents one of the primary challenges in our system.
+Nevertheless, our multi-stage pipeline design still demonstrates substantial overall gains, and the benefits significantly outweigh the drawbacks introduced by occasional decomposition errors.
+As shown in Table III, the performance has decreased when removing the editing (“w/o edit”) or re-retrieval (“w/o re-retrieval”) stage respectively, which demonstrates that ARE can address the error propagation.
+TABLE VI: The quality report of the generated instruct-tuning dataset
+   Dataset    Human    LLM-as-a-Judge
+   Train (500)    0.934    0.965
+   Test (200)    0.946    0.972
+   average    0.940    0.969
+We evaluated the quality of the generated instruct-tuning dataset (randomly selected 500 and 200 from train and test respectively) using both human evaluation and LLM-as-a-Judge (Llama-3.1-70B-Instruct).
+The accuracy scores for LLM generated long-form text, molecular clauses, and atomic facts were 0.94 (human) and 0.969 (LLM), as shown in Table VI.
+TABLE VII: Comparison of the proposed atomic fact-based retrieval with previous sub-question-based retrieval performance.
 Methods NQ StrategyQA ExpertQA
-$Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$ $Attr_{r}$ $Attr_{p}$
+$Attr_{r}$ $_{p} AttrrAtt$$_{r}$ $_{p} AttrrAtt$$_{r}$ $Attr_{p}$
 DRQA 0.424 0.647 0.237 0.49 0.127 0.283
@@
 […]
-5.4. The Impact of Fact Decomposition LLM with Different Iterations
-To find the best performance of molecular-to-atomic fact decomposition LLM and explore how iterations affect fact decomposition performance, we perform a comparative experiment, and the results are illustrated in Figure 4.
-Specificlly, we assess its performance at both the molecular and atomic levels, focusing on two key aspects: consistency and correctness.
-Consistency $n_{c}$ measures how closely the number of decomposed sentences aligns with that of gold sentences.
-Correctness $d_{correct}$ evaluates whether the decomposed sentences preserve the original sentence meaning.
-We test on a well-constructed evaluation dataset mentioned in 4.1 to select the decomposition model used in ARE.
-As shown in Figure 4, the performance of $n_{c}$ at the molecular-level initially increase and then decreasing as the number of epochs increase.
-At the atomic-level, $n_{c}$ shows an overall upward trend, stabilizing at epoch 4.
-For $d_{correct}$, performance at molecular-level also increase and then decrease as the number of epochs increase, peaking at epoch 3 with 0.972.
-However, at the atomic-level, it first decreases and then increases.
-In terms of consistency, the performance at the 3rd and 4th epochs is similar.
-However, for correctness, the score of the atomic-level at epoch 3 is the highest, and meanwhile the score of the molecular-level at epoch 3 and epoch 4 has little difference.
-Therefore, we select the LLM trained with 3 epochs.
-Details of evaluation process can be found in Appendix D.5.
-Table 4.
-Comparing the revised results with different systems.
+   Methods
+   Latency $\downarrow$
+   (sample/s)
+   Interaction $\downarrow$
+   (times/sample)
+   RARR    55    7.1
+   CCVER    87    9
+   ARE    41    4.7
+   Consumed tokens $\downarrow$
+   (tokens/sample)
+   Retrieval $\downarrow$
+   (times/sample)
+   RARR    7371    5.93
+   CCVER    9711    7
+   ARE    5340    4.77
+V-H Hypothesis Proof Experiment
+As illustrated in Table VII, ARE shows the improvements of 4.6% on NQ, 8.3% on StrategyQA, and 9.9% on ExpertQA based on the metric of $Attr_{r}$.
+This supports the hypothesis that atomic facts are easier to retrieve as relevant evidence.
+Furthermore, DRQA, using the entire answer and question for retrieval, struggles in cross-document retrieval scenarios.
+Methods involving generated sub-questions fail to obtain more evidence due to the poor quality of generated questions.
+CCVER and RARR both generate sub-questions and then use them for retrieval.
+Their performance is similar on NQ and ExpertQA, but on StrategyQA, CCVER outperforms RARR.
+This is because CCVER generates “Yes or No” type questions, which match the type of the StrategyQA dataset, allowing for better evidence retrieval.
+However, both perform worse than ARE, which is based on atomic fact retrieval.
+V-I Inference Efficiency Analysis
+To compare the inference efficiency between ARE and other baselines, we perform controlled inferences CCVER, RARR and ARE on 4 NVIDIA A800 GPUs using VLLM [^9] https://github.com/vllm-project/vllm.
+We measure average latency, interaction times with LLMs, and token consumption.
+As shown in Table VIII, ARE achieves a shorter average inference time per sample compared to RARR.
+The average interaction times with the LLMs are 4.7, and the average tokens consumed are 5340, both of which outperform the baselines.
+This is because the number of atomic facts generated by instruction-tuned decomposition LLM is less, but they are more relevant to the original answer than RARR and CCVER, leading to fewer retrievals (4.77 times for ARE, 5.93 for RARR and 7 for CCVER).
+TABLE IX: Comparing the revised results with different systems.
 Red indicates hallucinated content that conflicts from the evidence or fact.
@@
 ”Daddy’s Little Girl” is a song by American traditional pop and jazz singer Al Martino. The song was released in 2006 as the second single from her debut album ”Small Town Girl.”       Pres: 0.83
-5.5. Hypothesis Proof Experiment
-As illustrated in Table 3, ARE shows the improvements of 4.6% on NQ, 8.3% on StrategyQA, and 9.9% on ExpertQA based on the metric of $Attr_{r}$.
-This supports the hypothesis that atomic facts are easier to retrieve as relevant evidence.
-Furthermore, DRQA, using the entire answer and question for retrieval, struggles in cross-document retrieval scenarios.
-Methods involving generated sub-questions fail to obtain more evidence due to the poor quality of generated questions.
-CCVER and RARR both generate sub-questions and then use these sub-questions for retrieval.
-Their performance is similar on NQ and ExpertQA, but on StrategyQA, CCVER outperforms RARR.
-This is because CCVER generates “Yes or No” type questions, which match the type of the StrategyQA dataset, allowing for better evidence retrieval.
-However, both perform worse than ARE, which is based on atomic fact retrieval.
-5.6. Visualization results of attribution scores and preservation
-To provide a more intuitive comparison between existing methods and our proposed ARE, regarding the simultaneous achievement of maximum attribution scores and preservation of original intent, we conducted the visualization experiments illustrated in Figure 5.
-As shown in Figure 5, ARE is positioned on the right side of the dashed line, which indicates that ARE has achieved effective editing.
-The contour lines indicate that the proposed ARE achieves the best performance in both the F1RP (F1RP both considers Attrr and $Pres$) and F1PP (F1RP simultaneously considers Attrp and $Pres$) metrics after editing.
-These phenomena demonstrates the robustness of the proposed ARE in maximizing the attribution scores and preserving the original intent.
-The completed results can be found in Appendix D.4.
-5.7. Case Study
-When comparing the revised results of EFEC, RARR, and ARE in Table 4, clearly highlights its advantages, as well as the shortcomings of RARR and EFEC.
+TABLE X: Error case in ARE. Red indicates the hallucinated content. Blue marks as the correct content. Orange represents the invalid evidence. Green is the specialized word.
+Atomic Fact The Euro is the official currency of Spain and is used by 19 of the 27 European Union countries.
+case 1 Retrieved evidence …The euro is the official currency of 20 of the 27 member states of the European Union.
+Atomic Fact Chile is one of the countries with the highest forest cover in the world.
+case 2 Retrieved evidence This is far more than the second greatest country in terms of forest area coverage, which is Russia . Russia has approximately ….
+Long-form answer (Mintaka) The first book in the Hunger Games series, titled “The Hunger Games,” was published in 2008. (Total tokens: 16)
+Atomic Fact The first book was published in 2008.
+case 3 Retrieved evidence The Center for Fiction The Center for Fiction 2008 First Novel Prize …. This annual award was created in 2006 to honor the best first novel of the year.
+Long-form answer (ExpertQA) A client with substance addiction, ….
+Secondly, Dialectical Behavior ….
+Additionally, Cognitive-Behavioral Therapy (CBT) ….
+Finally, a trauma-informed approach to addiction….
+(Total tokens: 111)
+Atomic Fact A client with substance addiction, … requires a comprehensive and integrative therapeutic approach.
+case 4 Retrieved evidence … of childhood trauma in a range of settings that are particularly likely to serve this population like addictions, … in the context of the past trauma.
+V-J Case Study
+When comparing the revised results of EFEC, RARR, and ARE in Table IX, it clearly highlights its advantages, as well as the shortcomings of RARR and EFEC.
 ARE performs minimal edits by focusing on atomic facts, successfully preserving the original intent while improving factuality and attribution.
@@
 Such drastic revisions can result in a final product that significantly deviates from the original content and intent.
-6. Conclusion
+As shown in Table X, cases 1 and 2 illustrate verification errors where fine-grained differences between evidence and facts are missed, leading to incorrect classifications (e.g., labeling “editing required” as “supportive” or vice versa).
+We randomly selected 50 samples for manual evaluation, finding a 23.1% error rate in misclassifying “editing required” as “irrelevant”.
+In case 1, the evidence cites 20 Eurozone countries while the fact mentions 19, but this discrepancy is not detected.
+Case 2 presents irrelevant evidence about Russia’s forests for a fact concerning Chile.
+Case 3 shows a decomposition failure where the LLM fails to extract the key subject “The Hunger Games” from a long-form answer, resulting in an invalid atomic fact.
+Case 4 highlights compounded failure in medical-domain content involving a long (111-token) multi-fact structure and inductive reasoning.
+The decomposition is insufficient, leading to ineffective retrieval and cascading errors across subsequent stages.
+TABLE XI: The prompt templates used in ARE.
+LLMs
+Prompts
+Long-form
+answer
+generation
+You need to think step by step and answer my question.
+Example 1:
+[…]
+Please answer the question and give the Explanation and
+an answer.
+Evidence
+verifying
+I will check some things you said. Here are some
+examples for you to learn this process:
+Example 1:
+[…]
+[…]
+Now, please follow the above examples to inference.
+Fact
+expansion
+Rewrite the fact into two short atomic phrases based on
+Wikipedia. Ensure they contain more factual information
+and can be easily supported by search engines.
+Editing
+You need use some evidences to check ’You said’ if
+there are some difference and generate in ”My fix”.
+Here are some examples to learn:
+[…]
+[…]
+I will give you a new instance, please follow the above
+example to fix a new one and start with ”My fix:”.
+Backtracking
+You are a linguist. Your task is to compose coherent
+sentences using the provided atomic facts. While doing
+so, refer to the structure of the given reference text —
+including sentence count, subject order, and pronoun
+usage — to maintain consistency and fluency.
+Here are some examples:
+{ “reference answer”: long-form answer
+[{Molecular clause 1: [Atomic facts]},
+{Molecular clause 2: [Atomic facts]},
+{ Molecular clause 3: [Atomic facts] }]}
+Now, think step by step to merge the given facts by
+referring to the reference text. Generate the output in
+coherent sentences starting with ”Sentences:”.
+VI Conclusion
 This paper proposes a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for post-hoc retrieval-based AQA tasks, which contains a Molecular-to-Atomic decomposition stage and an Atomic-to-Molecular backtracking process.
@@
 The effectiveness of this framework is demonstrated across four datasets using four prominent LLMs, complemented by an extensive ablation study and LLM evaluation.
-Appendix
-Appendix A Prompt for LLMs
-A.1. Prompt for Long-form answer generation
-Prompt for Long-form answer generation you need think step by step and answer my question.
-Example 1: 1.
-Question: Which actor was the star of Titanic and was born in Los Angeles, California?
-2.
-Explanation: Titanic is a film directed by James Cameron.
-Leonardo DiCaprio was played the character Jack Dawson in Titanic.
-Leonardo DiCaprio was born in Los Angeles,California.
-3.
-Answer: Leonardo DiCaprio.
-Example 2: 1.
-Question: How many teams has Matthew Stafford played for?
-2.
-Explanation: Matthew Stafford is an American football quarterback.
-Matthew Stafford played for the Lions from 2009 to 2020.
-Matthew Stafford played for Los Angeles Rams in 2021.
-3.
-Answer: 2.
-Example 3: 1.
-Question: Is Jake Gyllenhaal older than Maggie Gyllenhaal?
-2.
-Explanation: Jake Gyllenhaal is an American actor.
-Jake Gyllenhaal was born in December 19, 1980, and Maggie Gyllenhaal is an American actress.
-Gyllenhaal was born in November 16, 1977.
-3.
-Answer: No.
-Example 4: 1.
-Question: How many sacks does Clay Matthews have in his career?
-2.
-Explanation: Clay Matthews has been pro linebacker for 11 seasons.
-Ten of the seasons Clay Matthews played for Green Bay Packers,while Clay Matthews last season was with Los Angeles Rams, dedicated to Clay Matthews hometown.
-Clay Matthews III had 383 solo tackles, 136 assists, and 91.5 sacks.
-Clay Matthews father, Clay Matthews Jr., also played in NFL and had 69.5 sacks in his career.
-3.
-Answer: 91.5 (Clay Matthews III) and 69.5 (Clay Matthews Jr.).
-please answer the Question: question and give the Explanation and a answer start with the ”Explanation” and ”Answer”.
-A.2. An example instruction data for instruction-tuning
-An example data for instruction-tuning {“instruction”: “Please extract the clauses and the corresponding atomic facts from the following text, when the pronoun appears, output the real name of the person or thing it refers to.”, input”: “Lawrence Wackett’s occupation is an aerospace engineer.
-He was involved in World War I and received the James Cook Medal on January 1, 1978.
-Lawrence Wackett’s gender is male.”, output”: [{“Lawrence Wackett’s occupation is an aerospace engineer.”: [“Lawrence Wackett’s occupation is an aerospace engineer.”]}, {’He was involved in World War I and received the James Cook Medal on January 1, 1978.’: [’Lawrence Wackett was involved in World War I.’, ’Lawrence Wackett received the James Cook Medal on January 1, 1978.’]}, {“Lawrence Wackett’s gender is male.”: [“Lawrence Wackett’s gender is male.”] }] }
-A.3. Prompt for evidence verifing and fact expansion
-Prompt for evidence verifing I will check some things you said.
-Here are some examples for you to learn this process: Example 1: 1.
-You said: Your nose switches back and forth between nostrils.
-When you sleep, you switch about every 45 minutes.
-This is to prevent a buildup of mucus.
-It’s called the nasal cycle.
-2.
-I checked: When you sleep, you switch about every 45 minutes.
-3.
-I found this article: Although we don’t usually notice it, during the nasal cycle one nostril becomes congested and thus contributes less to airflow, while the other becomes decongested.
-On average, the congestion pattern switches about every 2 hours, according to a small 2016 study published in the journal PLOS One.
-4.
-Reasoning: The article said the nose’s switching time is about every 2 hours, and you said the nose’s switching time is about every 45 minutes.
-5.
-Therefore: This disagrees with what you said.
-Example 2: 1.
-You said: The Little House books were written by Laura Ingalls Wilder.
-The books were published by HarperCollins.
-2.
-I checked: The Little House books were published by HarperCollins.
-3.
-I found this article: These are the books that started it all – the stories that captured the hearts and imaginations of children and young adults worldwide.
-Written by Laura Ingalls Wilder and published by HarperCollins, these beloved books remain a favorite to this day.
-4.
-Reasoning: The article said the Little House books were published by HarperCollins and you said the books were published by HarperCollins.
-5.
-Therefore: This agrees with what you said.
-Example 3: 1.
-You said: Real Chance of Love was an American reality TV show.
-Season 2 of the show was won by Cali, who chose to be with Chance.
-2.
-I checked: Season 2 of the show was won by Cali.
-3.
-I found this article: Real Chance of Love 2: Back in the Saddle is the second season of the VH1 reality television dating series Real Chance of Love.
-Ahmad Givens (Real) and Kamal Givens (Chance), former contestants on I Love New York are the central figures.
-4.
-Reasoning: The article doesn’t answer the question and you said that Cali won season 2 of Real Chance of Love.
-5.
-Therefore: This is irrelevant to what you said.
-Example 4: 1.
-You said: Tiger Woods is the only player who has won the most green jackets.
-He has won four times.
-The Green Jacket is one of the most coveted prizes in all of golf.
-2.
-I checked: The Green Jacket is one of the most coveted prizes in all of golf.
-3.
-I found this article: The green jacket is a classic, three-button, single-breasted and single-vent, featuring the Augusta National Golf Club logo on the left chest pocket.
-The logo also appears on the brass buttons.
-4.
-Reasoning: The article said the Green Jacket is a classic three-button single-breasted and single-vent and you said the Green Jacket is one of the most coveted prizes in all of golf.
-5.
-Therefore: This is irrelevant to what you said.
-Now, please follow above examples pattern to inference new one (don’t output the examples), and outputs the result start with ’4.
-Reasoning: ’ and ’5.
-Therefore: ’.
-If from the article supports about ’you said’, you should generate ”agrees”, if unsupports, you shuld generate ”disagrees” and if the article is not fully support ’you said’, you should generate ”irrelevant” in ’Therefore’.
-1.
-You said: {claim 2.
-I checked: {query} 3.
-I found this article: {evidence}
-Prompt for fact expansion Rewrite the fact: {fact} into two short atomic phrases based on Wikipedia. Ensure they contain more factual information and can be easily supported by search engines.
-A.4. Prompt for editing
-Prompt for atomic editing you need use some evidences to check ’You said’ if there are some difference and generate in ”My fix”.
-Here are some examples to learn: Example 1: 1.
-You said: 11 Diagonal Street was actually completed way back in 1978.
-2.
-I found these evidences: 11 Diagonal Street completion date ”1983” .
-11 Diagonal Street building end date ”1983” .
-11 Diagonal Street building start date ”1978”.
-3.
-This suggests 11 Diagonal Street was completed in 1978 is wrong.
-4.
-My fix: 11 Diagonal Street was actually completed way back in 1983.
-Example 2: 1.
-You said: The length of Mississippi River is 2340 km.
-2.
-I found these evidences: Missouri River length +3726.
-Mississippi River discharge +12743.
-Mississippi River tributary Yazoo.
-3.
-The question is What is the length of the Mississippi River?
-This suggests 2340 km of Mississippi River in your statement is wrong.
-4.
-My fix: The length of Mississippi River is 3726 km.
-Example 3: 1.
-You said: Your nose switches back and forth between nostrils.
-When you sleep, you switch about every 45 minutes.
-This is to prevent a buildup of mucus.
-It’s called the nasal cycle.
-2.
-I found this evidences: Although we don’t usually notice it, during the nasal cycle one nostril becomes congested and thus contributes less to airflow, while the other becomes decongested.
-On average, the congestion pattern switches about every 2 hours, according to a small 2016 study published in the journal PLOS One.
-3.
-This suggests 45 minutes switch time in your statement is wrong.
-4.
-My fix: Your nose switches back and forth between nostrils.
-When you sleep, you switch about every 2 hours.
-This is to prevent a buildup of mucus.
-It’s called the nasal cycle.
-Example 4: 1.
-You said: In the battles of Lexington and Concord, the British side was led by General Thomas Hall.
-2.
-I found this evidences: Interesting Facts about the Battles of Lexington and Concord.
-The British were led by Lieutenant Colonel Francis Smith.
-There were 700 British regulars.
-3.
-This suggests General Thomas Hall in your statement is wrong.
-4.
-My fix: In the battles of Lexington and Concord, the British side was led by Lieutenant Colonel Francis Smith.
-Now, I will give you a new instance, please follow the above example to fix new one and start with ”My fix:”, do not generate irrelevant information.
-1.
-You said: {claim} 2.
-I found these evidences: {evidence} 3.
-This suggests
-Appendix B Theoretical Analysis of the Fact Expansion
-When the retrieved evidence is invalid, we consider the possible reasons to be: The informativeness of the query contained may too lower.
-Increasing the informativeness of the query is an avenue that can be tried, and and the process will be justified in terms of informativeness.
-The informativeness of an atomic fact can be defined as:
-$I(AF_{ij})$ $=-\log P(\boldsymbol{AF}_{ij})$
-$=P(t_{1},\dots,t_{i},\dots,t_{n})$
-$=P(t_{1})\cdot\dots\cdot P(t_{n}\mid t_{1},t_{2},\dots,t_{n-1}),$
-where $t_{i}$ is a token in $AF_{ij}$.
-We assume that the set of facts that can be extended from $AF_{ij}$ is $S_{ex}=\{ef_{1},...,ef_{j},...,ef_{n}\}$, then the information content of each $ef_{j}$ is $I(ef_{j})$.
-Finally, for an invalid query, we expand it to $AF_{ij}^{ex}$, then its final information content is
-$I(AF_{ij}^{ex})$ $=-\log P(AF_{ij})-\log P(ef_{i})$
-$=I(AF_{if})+I(ef_{j})>I(AF_{if}).$
-Table 5. The experimental results of aggregating evidence retrieved from sub-questions into the corresponding molecular clauses.
-GPT-3.5 NQ StrategyQA ExpertQA
-[…]
-RARR 0.649 0.058 0.868 0.109 0.743 0.356 0.097 0.862 0.174 0.504 0.34 0.078 0.904 0.144 0.494
-RARR w aggregate evidnece 0.499 0.646 0.942 0.766 0.653 0.336 0.53 0.948 0.680 0.496 0.276 0.407 0.959 0.571 0.428
-ARE 0.67 0.756 0.91 0.826 0.772 0.463 0.559 0.899 0.689 0.611 0.412 0.438 0.917 0.593 0.569
-[…]
-Appendix C The details of the Experimental Settings
-C.1. Datasets
-Natural Questions is a question-answering dataset developed by the Google search engine and is widely used to evaluate machine reading comprehension, information retrieval, and other tasks.
-Its questions are complex and often ambiguous, requiring the identification of relevant answers from long documents, making it a knowledge-intensive task.
-Mintaka is a complex multilingual dataset featuring superlative, cross, and multi-hop question types.
-It is based on knowledge graphs (KG), which are also knowledge-intensive datasets widely used to evaluate end-to-end deep learning question-answering models.
-StrategyQA is a dataset focused on open-domain problems, which involve inference steps in the questions and require the model to provide the reasoning process to arrive at the correct answer.
-This places a high demand on the model’s attribution ability, as its answers typically reference different paragraphs from multiple documents.
-ExpertQA is a dataset specifically designed to evaluate attribution-based question answering models.
-It encompasses seven question types across 32 domains, and answering these questions requires specialized knowledge, posing a greater challenge for attribution systems.
-C.2. Baselines
-  • DRQA [8] is a straightforward attribution model that concatenates the question and answer into a single query before performing retrieval.
-The purpose of this baseline is to demonstrate how much data in the dataset does not require decomposition prior to retrieval.
-  • RARR [9] is the first method to implement a retrieval-then-revise paradigm in the AQA task, with all its components based on API-driven LLMs.
-It first prompts the LLM to generate a long-form answer as a claim, then generates a series of sub-questions based on this claim, using these sub-questions as queries to retrieve evidence from a search engine.
-RARR has also shown that using Bing Search and Google Search as evidence sources yields similar performance.
-  • EFEC [41] is an editor built on the T5 model that modifies a claim by incorporating evidence.
-It is trained using both weak and full supervision methods.
-To demonstrate its optimal performance, we reproduced it using full supervision.
-EFEC employs a deep neural network-based retriever for evidence retrieval.
-For a fairer comparison, we adopted the same settings as RARR, generating sub-questions for retrieval and using the retrieved evidence to edit the answer.
-  • CCVER [42] is a model focused on fact verification, consisting of a two-stage retrieval process.
-In the first stage, it uses a few-shot approach to prompt LLMs to generate Yes or No-type sub-questions for a given claim, which are then used as queries to retrieve relevant evidence.
-In the second stage, it combines BM25 and LLMs to summarize the evidence.
-Finally, DeBERTa-large is employed to evaluate the claim based on the evidence.
-We selected CCVER to explore the impact of generating different types of questions on attribution performance, utilizing only its question generation prompt in conjunction with RARR’s editing method.
-Since the models reported in the RARR paper (GPT-3 and PaLM) are no longer available, we used other LLMs to reproduce RARR.
-Table 6. Ablation experimental results of GPT3.5 on Mintaka
-Methods Mintaka
-[…]
-w/o edit 0.672 0.768 - - -
-w/o atomic 0.477 0.651 0.928 0.765 0.630
-w/o molecular 0.725 0.832 0.548 0.661 0.624
-w/o re-retrieval 0.680 0.767 0.936 0.843 0.788
-ARE 0.716 0.807 0.914 0.857 0.803
-C.3. Metric
-NLI model.
-We use an NLI model based on T5-11B, which is trained on six datasets: MNLI, SNLI, FEVER, PAWS, SciTail, and VitaminC  [43, 44, 45, 46, 47, 48].
-The input format is: “premise: {evidence} hypothesis: {sentences}”.
-As suggested by [9], we use the probability of producing “1” as the entailment score in the metric $Attr_{r}$ and use the binary result “1” (entailed) or “0” (not entailed) for $Attr_{p}$.
-Levenshtein distance.
-Preservation generally utilizes Levenshtein distance to measure the changed information from $X$ to $X^{\prime}$, which assesses the difference between two strings.
-Levenshtein distance is defined as the minimum number of single-character edits (insertions, deletions, or substitutions) required to transform one string into another.
-This metric quantifies the dissimilarity between two sequences and is commonly used in applications such as spell-checking and natural language processing.
-C.4. Other Settings in Experiments.
-The Complete Heuristics Rules
-The complete heuristics rules used in instruction-tuning dataset construction Triples containing the following will be filtered: “ID”, “.svg”, “.jpg”, “.png”, “.JPG”, “.JPEG”, “http://”, “https://”, “Category:”, “Wikipedia:”, “code”,“UMLS CUI”, “Wikimedia”, “.map”,“ISO”,“code”, “UNESCO”,“html” , “IPTC” ,“ISNI” , “ISSN”.
-Molecular-to-Atomic fact decomposition LLM.
-To instruction-tune the Molecular-to-Atomic fact decomposition LLM, we used Llama-Factory [^9] https://github.com/hiyouga/LLaMA-Factory [49] for fine-tuning.
-The hyperparameters used in the training are: The number of epochs is 3, learning rate is 1e-4, batch_size is 1, warmup_ratio is 0.1, cutoff_len is 2048, per_device _train_batch_size is 1 and gradient_accumulation_steps is 2, lr _scheduler_type is cosine.
-The entire training and inference are conducted on 2 * A800 80G GPUs.
-LLMs.
-To efficiently infer the LLaMA series of large language models, we utilized VLLM [^10] https://github.com/vllm-project/vllm for inference on two A800 GPUs.
-To ensure reproducibility and stability, all temperature settings are set to 0, except for the Llama3 series, for which we adjusted the temperature to the lowest possible value of 0.1, as it cannot be set to 0.
-Table 7. Ablation experiment results based on GPT-4o-mini
-[…]
-GPT-4o-mini NQ StragegyQA
-w/o edit 0.587 0.706 - - - 0.364 0.467 - - -
-w/o atomc 0.328 0.541 0.884 0.671 0.479 0.189 0.237 0.883 0.374 0.280
-w/o molecular 0.653 0.72 0.544 0.620 0.594 0.429 0.514 0.563 0.537 0.487
-w/o re-retrieval 0.602 0.691 0.924 0.791 0.729 0.402 0.462 0.914 0.614 0.558
-ARE 0.623 0.722 0.902 0.802 0.737 0.413 0.469 0.904 0.618 0.567
-Mintaka ExpertQA
-w/o edit 0.634 0.766 - - - 0.377 0.401 - - -
-w/o atomc 0.349 0.566 0.870 0.686 0.499 0.195 0.290 0.921 0.441 0.321
-w/o molecular 0.648 0.76 0.446 0.562 0.528 0.391 0.415 0.549 0.473 0.457
-w/o re-retrieval 0.619 0.707 0.871 0.780 0.723 0.369 0.390 0.939 0.551 0.529
-ARE 0.665 0.755 0.881 0.813 0.738 0.387 0.414 0.914 0.570 0.545
-Appendix D Other experimental results
-D.1. The Ablation Study on Mintaka with GPT-3.5
-As shown in Table 6, similar to the other datasets, each module in ARE enhances attribution, particularly the editing module, which increases $Attr_{r}$ by 0.044 and $Attr_{p}$ by 0.039 on the Mintaka dataset.
-D.2. Experimental results on GPT-4o-mini.
-We also conducted experiments on the latest GPT series model, GPT-4o-mini (To ensure reproducibility, we used GPT-4o-mini-2024-07-18.). The experimental results can be found in Table 8.
-We also conducted ablation experiments on GPT-4o-mini, and the results can be found in Table 7.
-The experimental results show that ARE also achieved the best performance on GPT-4o-mini, maximizing the attribution score while preserving the original intent.
-Table 8. Experimental results using GPT-4o-mini on four datasets.
-GPT-4o-mini
-[…]
-NQ
-DRQA 0.265 0.576 - - -
-EFEC 0.452 0.053 0.688 0.098 0.546
-CCVER 0.563 0.051 0.931 0.097 0.702
-RARR 0.493 0.08 0.904 0.147 0.638
-ARE 0.623 0.722 0.902 0.802 0.737
-Mintaka
-DRQA 0.251 0.544 - - -
-EFEC 0.497 0.049 0.698 0.092 0.581
-CCVER 0.634 0.045 0.933 0.086 0.755
-RARR 0.508 0.058 0.899 0.109 0.649
-ARE 0.634 0.755 0.881 0.813 0.737
-StrategyQA
-DRQA 0.08 0.222 - - -
-EFEC 0.33 0.071 0.638 0.128 0.435
-CCVER 0.341 0.056 0.961 0.106 0.503
-RARR 0.292 0.079 0.916 0.145 0.443
-ARE 0.413 0.469 0.904 0.618 0.567
-ExpertQA
-DRQA 0.122 0.308 - - -
-EFEC 0.297 0.075 0.635 0.134 0.405
-CCVER 0.31 0.062 0.98 0.117 0.471
-RARR 0.269 0.064 0.944 0.120 0.419
-ARE 0.387 0.414 0.914 0.570 0.544
-D.3. Experimental results of aggregating evidence from sub-questions.
-The sub-question generation method (such as RARR) lacks a clear correspondence between questions and clauses, failing to provide complete evidence for molecular clauses.
-Therefore, we employ our Molecular-to-Atomic fact decomposition LLM to first decompose the long-form answer.
-Next, based on the molecular clauses, we utilize the sub-question generation method from RARR to prompt LLMs to generate sub-questions for retrieval.
-Finally, the evidence retrieved through the sub-questions is aggregated according to its relationship with the molecular clauses, serving as supporting evidence for these clauses.
-As shown in Table 5, when evidence was aggregated for molecular clauses, the $Attr_{p}$ score improved significantly, increasing by 0.588 on NQ, 0.433 on StrategyQA, and 0.329 on ExpertQA.
-However, the $Attr_{r}$ metric is lower than that of RARR.
-This is because considering each molecular clause individually can lead to a loss of context, especially crucial subjects, resulting in unclear sub-questions that fail to retrieve key information.
-D.4. Visualization results of attribution scores and Pres.
-The visualization results for Mintaka and ExpertQA are presented in Figure 6.
-These figures illustrate the distribution of performance metrics, showing how different methods, including ARE, perform in terms of attribution and preservation of intent across these two datasets.
-D.5. The Evaluation process of Fact Decomposition LLM
-we define two metrics: $n_{c}$ for consistency and $d_{correct}$ for correctness.
-For consistency, it checks whether the number of decomposed sentences $s_{n}$ matches the number of gold sentences $g_{n}$.
-The decomposition is considered consistent when the number of decomposed sentences and gold sentences are equal.
-Formally, $n_{c}$ is defined as:
-$n_{c}=\mathbb{I}(\frac{s_{n}}{g_{n}}=1),$
-where $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if the condition is satisfied (i.e., the ratio equals 1) and 0 otherwise.
-This metric ensures that the decomposition does not produce too few or too many sentences relative to the gold reference.
-For correctness, it is evaluated whether the meaning of the decomposed sentences remains consistent with the gold sentences.
-Let $d_{s}$ represent a decomposed sentence and $g_{s}$ a gold sentence.
-The correctness assessed using a binary NLI result with TRUE model, to check if $d_{s}$ entails $g_{s}$ and vice versa, it can be defined as:
-$d_{correct}=\frac{NLI_{bi}(d_{s},g_{s})}{s_{n}}.$
-Here, $NLI_{bi}(d_{s},g_{s})$ measures the entailment between decomposed and gold sentences.
-The score is normalized by the total number of decomposed sentences $s_{n}$ , ensuring that we account for the number of decomposed sentences when calculating correctness.
-This metric ensures that each decomposed sentence accurately reflects its corresponding gold sentence in terms of meaning.
-Appendix E The Proof of $Attr_{p}$ can be used to Evaluate the Precision
-The Precision can be defined as:
-Precision=True Positive (TP)True Positive (TP)+False Positive ($\text{Precision}=\frac{\text{True Positive (TP)}}{\text{True Positive (TP)}+\text{False Positive (FP)}}.$
-In this paper, “True Positive (TP)” is $E_{i}$ can correctly entail at least one molecular clause $MF_{i}$, “False Positive (FP)” indicates $E_{i}$ do not entail any of the molecular clauses $MF_{i}$.
-According to the equation 6, $\text{NLI}_{bi}(E_{i},MF_{i})=1$ if $E_{i}$ entails $MF_{j}$, and otherwise.
-for each $E_{i}$, the number of FP can be calculated:
-$\text{FP}=\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=0),$
-For TP, it can be calculated as:
-$\text{TP}=\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=1),$
-The attribution report collection $\{E_{1},...,E_{m}\}$ has $m$ evidence, each of them considered as a predicted positive instance. Therefore, the total number of predicted positives is:
-Total Predicted $\text{Total Predicted Positives}=\text{TP}+\text{FP}=m.$
-The number of TP is: $\text{TP}=m-\text{FP}$. Applying the precision formula:
-Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}}$
-$=\frac{\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i})=1)}{m},$
-due $\text{NLI}_{bi}$ will return the binary result, it can be further simplified to:
-Precision $=\frac{\sum_{i=1}^{m}\mathbb{I}(\mathrm{NLI}_{bi}(E_{i},MF_{i}))}{m}$
-$=Attr_{p}.$
+VII ACKNOWLEDGMENT
+This research was supported by NSFC (No.62376144), by Science and Technology Cooperation and Exchange Special Project of Shanxi Province (No.202204041101016), by Key Research and Development Project of Shanxi Province (No.202102020101008), by Natural Language Processing Innovation Team (Sanjin Talents) Project of Shanxi Province.
+Jiaoyan is funded by the EPSRC project OntoEm (EP/Y017706/1)